### v2

# Bandit Sequential Posted Pricing via Half-Concavity

Abstract
Sequential posted pricing auctions are popular because of their simplicity in practice and their tractability in theory. A usual assumption in their study is that the Bayesian prior distributions of the buyers are known to the seller, while in reality these priors can only be accessed from historical data. To overcome this assumption, we study sequential posted pricing in the bandit learning model, where the seller interacts with $n$ buyers over $T$ rounds: In each round the seller posts $n$ prices for the $n$ buyers and the first buyer with a valuation higher than the price takes the item. The only feedback that the seller receives in each round is the revenue.
Our main results obtain nearly-optimal regret bounds for single-item sequential posted pricing in the bandit learning model. In particular, we achieve an $\widetilde{O}(\mathsf{poly}(n)\sqrt{T})$ regret for buyers with (Myerson’s) regular distributions and an $\widetilde{O}(\mathsf{poly}(n)T^{{2}/{3}})$ regret for buyers with general distributions, both of which are tight in the number of rounds $T$. Our result for regular distributions was previously not known even for the single-buyer setting and relies on a new half-concavity property of the revenue function in the value space. For $n$ sequential buyers, our technique is to run a generalized single-buyer algorithm for all the buyers and to carefully bound the regret from the sub-optimal pricing of the suffix buyers.
1 Introduction
Sequential Posted Pricing (SPP) schemes are well-studied in mechanism design because of their simplicity/convenience in practice and their tractability in theory. In the basic setting, a seller wants to sell a single item to a group of buyers. The buyers arrive one-by-one and the seller presents a take-it-or-leave-it price. The first buyer with a value higher than the posted price takes the item by paying the price. The benefit of SPP is that it gives approximately-optimal revenue while being “simple”. For instance, although it is known that Myerson’s mechanism [Mye81] is optimal for selling a single item to $n$ buyers with regular distributions, this mechanism is often impractical. This is because the winner’s payment is defined via “complicated” virtual valuations and all buyer bids need to be simultaneously revealed, which is not possible in large markets. In contrast, SPP are known to give at least $1/2$-fraction of the optimal revenue for regular buyers, while having fixed prices and buyers arriving one-by-one (see the books [Rou17, Har13] for regular buyers, and [Yan11] for a discussion on how ironing extends this result to general buyers). For multi-item (multi-parameters) setting, the difference between optimal and SPP mechanisms is even more stark. For instance, the optimal mechanism to sell $n$ items to a single unit-demand bidder is known to be impossible (unless $P^{NP}=P^{\#P}$) [CDO+15], but we can use SPP to obtain $1/4$-fraction of the optimal revenue [CHK07, CHMS10, CMS15].
Sample Complexity. A common assumption among initial works on revenue-maximization was that the underlying distributions of the buyers are known to the seller. This is unrealistic in many applications since the distributions are unknown and need to be learnt from historical data. Inspired by this, a recent line of research studies the sample complexity for different types of mechanisms, i.e., how many samples are sufficient to learn an $\epsilon$-optimal mechanism. For instance, starting with the pioneering work in [CR14, BBHM08], several papers studied the sample complexity of Myerson’s mechanism [HMR18, DHP16, RS16, GN17], and finally [GHZ19] obtained the tight sample complexity bounds for regular buyers and for $[0,1]$ bounded-support buyers. Sample complexity of many other classes of mechanisms have been studied, e.g., [BSV18, JLX23] study second-price auctions, [BSV18, BDD+21] study posted-pricing mechanisms, [GHTZ21] study strongly-monotone auctions (which includes Myerson and SPP), and [MR16, CD17, GW21, COZ22] study the sample complexity of multi-parameter auctions.
Bandit Learning. A stronger model of learning than the sample complexity model is the well-studied bandit learning model; see books [LS20, BC12]. In this model, the seller interacts with the buyers over $T$ days. On each day, the seller proposes a parameterized mechanism to the buyer and sees only the revenue as feedback. The goal is to minimize the total regret, which is the difference between the total revenue achieved by the learning algorithm and the optimal mechanism over $T$ days. It is stronger than sample complexity in two aspects: Firstly, it requires the learner to continually play near-optimal mechanisms, whereas in sample complexity the learner may lose a lot of the revenue while learning. Secondly, the feedback in the bandit model on any day is limited to only the proposed mechanism, and not every possible mechanism that could have been proposed.
The study of learning mechanisms with bandit feedback goes back to at least [KL03], where the authors provide an $\widetilde{O}(T^{{2}/{3}})$ regret bound for learning single-buyer posted pricing mechanism. Other examples of learning auctions in this model include bandits with knapsacks [BKS18, ISSS22], reserve price for i.i.d buyers [CGM15], and non-anonymous reserve prices [NGW+21].
Despite a lot of work on learning SPP mechanisms and on learning auctions with bandit feedback, learning SPP in the bandit model was unknown. We answer this question and provide near-optimal regret bounds for SPP with $n$ buyers having regular or general distributions.
1.1 Model and Results
In Sequential Posted Pricing (SPP), there are $n$ sequential buyers with independent valuation distributions $\mathcal{D}_{1},\cdots,\mathcal{D}_{n}$. We make the standard normalization assumption that each distribution $\mathcal{D}_{i}$ is supported in $[0,1]$. The seller plays a group of prices $(p_{1},\cdots,p_{n})\in[0,1]^{n}$ and then the buyers arrive one-by-one with valuation $v_{i}\sim\mathcal{D}_{i}$. The first buyer with $v_{i}\geq p_{i}$ takes the item and pays the price $p_{i}$ as the revenue. The goal of the seller is to play prices to maximize the expected revenue. We define $R(p_{1},\cdots,p_{n})$ to be the expected revenue when playing prices $(p_{1},\cdots,p_{n})$, i.e.,
$R(p_{1},\cdots,p_{n})~{}:=~{}\sum_{i=1}^{n}p_{i}\cdot\mbox{\rm\bf Pr}\left[i\text{ gets the item}\right]~{}=~{}\sum_{i=1}^{n}p_{i}\cdot(1-F_{i}(p_{i}))\prod_{j=1}^{i-1}F_{j}(p_{j}),$
where $F_{i}(x)$ is the cumulative distribution function (CDF) of distribution $\mathcal{D}_{i}$. For SPP, the optimal prices are defined as
$(p^{*}_{1},\cdots,p^{*}_{n})~{}:=~{}{\arg\!\max}_{(p_{1},\cdots,p_{n})}R(p_{1},\cdots,p_{n}).$
These optimal prices can be easily calculated in polynomial time using a reverse dynamic program: price $p^{*}_{n}=\arg\!\max_{p}p(1-F_{n}(p))$, and with known optimal prices $p^{*}_{i+1},\cdots,p^{*}_{n}$, we can calculate
$p^{*}_{i}~{}=~{}\arg\!\max_{p}p(1-F_{i}(p))+F_{i}(p)\cdot\mbox{\rm\bf E}\left[\text{revenue from }i+1,\cdots,n\right].$
Bandit Sequential Posted Pricing (BSPP). In many practical applications of SPP, the valuation distributions $\mathcal{D}_{i}$ are unknown, so we will study them in the bandit learning model. Consider the following toy example to motivate the model: Suppose you want to sell an item each day for the entire next month using SPP. For simplicity, assume that each day exactly 3 bidders arrive: one each in the morning, afternoon, and evening. If you know the value distributions of these 3 bidders on each day, then you can find the optimal prices $(p^{*}_{1},p^{*}_{2},p^{*}_{3})$ as discussed above, and play them every day to maximize the total revenue. However, if the distributions are unknown then you need to learn the optimal prices during the month. A major challenge is that you don’t even get to see the true valuations of the 3 bidders who show up each day, only whether they decide to buy the item, or not, at the played price.
Formally, BSPP is a $T$ rounds/days repeated game where on each day $t\in[T]$ the seller proposes a group of $n$ prices $(p_{1}^{(t)},\cdots,p_{n}^{(t)})$. The buyers then arrive one-by-one with valuations $v_{i}^{(t)}\sim\mathcal{D}_{i}$ and the first buyer $i$ with valuation $v_{i}^{(t)}\geq p_{i}^{(t)}$ takes the item. The seller receives revenue $\text{Rev}^{(t)}=p_{i}^{(t)}$ as the reward for this day and sees $p_{i}^{(t)}$ as the feedback, or equivalently sees the identity of the buyer that takes the item [^1] By adding an arbitrarily small noise to each price, the seller can retrieve buyer’s identity from the revenue.. Note that none of the buyer valuations are ever revealed. If there is no buyer with a valuation higher than their price, the seller receives reward 00. The goal of the seller is to minimize in expectation the total regret:
$\text{Regret}~{}:=~{}T\cdot R(p^{*}_{1},\cdots,p^{*}_{n})-\sum_{t\in[T]}\text{Rev}^{(t)}.$
Our first main result gives a near-optimal regret algorithm for BSPP with regular buyers.
Main Result 1 (Informal 3.1).
For bandit sequential posted pricing with $n$ buyers having regular distribution inside $[0,1]$, there exists an algorithm with $\widetilde{O}(\mathsf{poly}(n)\sqrt{T})$ regret.
It should be noted that this is the first bandit learning algorithm for regular distributions with optimal regret bound in $T$, even for the single-buyer or i.i.d. buyers setting. A very recent paper of [LSTW23b] studies “pricing query complexity” of this problem in the special case of a single-buyer. Their model is weaker than our bandit learning model since it only gives a final regret bound, i.e., one might incur a large regret during the $T$ days of learning but the final learnt prices have a low $1$-day regret. Hence, their $\Omega({1}/{\epsilon^{2}})$ lower bound for single-buyer query complexity immediately implies an $\Omega(\sqrt{T})$ lower bound for BSPP, showing tightness of the $\sqrt{T}$ factor in 1. However, their single-buyer pricing query complexity upper bounds do not apply to our stronger bandit learning model. Moreover, it’s unclear how to extend their techniques beyond a single-buyer.
Our next result gives a near-optimal regret bound for BSPP with $n$ buyers having general distributions. The authors of [LSTW23b] give an $\Omega({1}/{\epsilon^{3}})$ query complexity lower bound for single-buyer setting with general distributions, implying an $\Omega(T^{2/3})$ regret lower bound for BSPP with general distributions. Our result achieves a tight upper bound dependency on $T$.
Main Result 2 (Informal 4.1).
For bandit sequential posted pricing problem with $n$ buyers having value distribution inside $[0,1]$, there exists an algorithm with $\widetilde{O}(\mathsf{poly}(n)T^{{2}/{3}})$ regret.
This result generalizes the single-buyer result of [KL03] to $n$ sequential buyers. Interestingly, our techniques are very different from theirs since we are in the stochastic bandit model. Although [KL03] show an $\widetilde{O}(T^{{2}/{3}})$ regret algorithm even when the buyer valuations are chosen by an adversary, we observe in Section 5 that such a result is impossible for multiple buyers since already for $2$ buyers with adversarial valuations, every online algorithm incurs $\Omega(T)$ regret.
1.2 Techniques
Single Regular Buyer via Half-Concavity. The proof of our first main result is based on a key observation that the revenue curve of a regular distribution is “half-concave”. Recall that a distribution is regular if its revenue curve is concave in the quantile space. Simple examples show that regular distributions need not be concave in the value space (e.g., exponential distributions). Our half-concavity shows that regular distributions are still concave on one side of its maximum.
Main Result 3 (Informal 2.4 and 3.9).
Let $R(p)$ be the revenue function of a regular distribution supported on $[0,1]$ and let $p^{*}=\arg\!\max_{p}R(p)$. Then, $R(p)$ is concave in $[0,p^{*}]$.
Concavity is a strong property that often allows efficient learning. We show that it’s sufficient to learn a single-peaked function even if it is only half-concave. The high-level intuition for learning a half-concave follows from the standard recursive algorithm for learning concave functions: first, consider a (fully) concave function $R(x)$ defined on interval $[\ell,r]$. We set $a=\frac{2\ell+r}{3}$ and $b=\frac{\ell+2r}{3}$, and test $R(a)$ and $R(b)$ with sufficiently many samples. There can be three cases:
  • Case 1: $R(a)<R(b)$. We can drop $[\ell,a]$ and recurse on $[a,r]$.
  • Case 2: $R(a)>R(b)$. We can drop $[b,r]$ and recurse on $[\ell,b]$.
  • Case 3: $R(a)\approx R(b)$. Concavity implies that $R(x)$ is nearly a constant for all $x\in[\ell,r]$, so we are done.
Now suppose $R(x)$ is only half-concave. We can still perform a similar algorithm: the first two cases use single-peakedness, so they remain the same. For the third case, half-concavity still guarantees that $R(x)$ can’t be too high for $x\in[b,r]$, so we can drop $[b,r]$ and recurse on $[\ell,b]$.
Our final algorithm and proofs combine the above Case 2 and 3; see details in Section 2.
Generalizing Single-Buyer to Sequential Buyers. In the case of sequential buyers, our main idea is to run the single-buyer algorithm for all $n$ buyers. A major difference compared to the single buyer setting is that the revenue function for this buyer becomes $R(p)=p\cdot(1-F(p))+C\cdot F(p)$, where $C\cdot F(p)$ represents the case that this buyer does not take the item and the seller receives an expected revenue $C$ from the buyers after the current buyer. Therefore, the main for the generalization is to take care of this extra $C\cdot F(p)$ term. We show that for regular distributions the half-concavity still holds for the new function $R(p)$ in the interval $[C,1]$, along with some other nice properties in the interval $[0,C]$. For general distributions, we show that the extra $C\cdot F(p)$ term does not make a huge difference.
1.3 Further Related Work
Bandit Learning and Single-Buyer Posted Pricing. The bandit learning model is well-established and we refer the readers to these books for classical results [LS20, BC12, CBL06]. The recent book of [Sli19] is a great reference for work at the intersection of bandits and economics. In particular, the problem of learning the single-buyer posted pricing mechanism with bandit feedback has a long history, dating back to [KL03]. In their paper, the authors put forth an $\widetilde{O}(T^{2/3})$ regret bound for general buyer distributions and an $\widetilde{O}(\sqrt{T})$ regret bound under a non-standard assumption that the revenue curve is single-peaked, and its second derivative at this peak is a strictly negative constant, independent of the parameter $T$. This result is not directly comparable to our $\widetilde{O}(\sqrt{T})$ bound for regular distributions. This is primarily because the second derivative at the maximum of the revenue curve for a regular distribution can range from negligibly small to zero. Additionally, [CBCP19] proposes an $\widetilde{O}(\sqrt{KT})$ regret bound, assuming the buyer’s value resides within a discrete set with cardinality $K$. This result, however, is also not directly comparable to ours, as regular distributions inherently exhibit continuity, contrasting with the assumption of discrete value sets.
Threshold Query Model. Recently, the threshold query structure that underpins the bandit feedback model of the single-buyer posted pricing problem is extensively studied. In this model, the learner queries a threshold $\tau$ and only observes $\mathbf{1}[\tau<X\sim\mathcal{D}]$. The goal is to determine the minimum number of queries to learn a key parameter (e.g., median, mean, or CDF) of a distribution. We refer the readers to [MN21, LSTW23b, OGKG23, LSTW23a] for learning complexity of the threshold query model.
Sequential Posted Pricing and Prophet Inequality. Sequential posted pricing (SPP) and its variants have been long popular, both in theory and in practice. One of the first results in their theoretical study is that posted prices obtain 78% of the optimal revenue for selling a single item in a large market [BH08]. Subsequently, posted prices have been successfully analyzed in both single-dimensional and multi-dimensional settings; see this survey [Luc17] and book [Har13]. A beautiful paper on the popularity of posted prices in practice is [EFLS18].
A closely related problem to SPP is the Prophet Inequality (PI) problem from optimal stopping theory. In both these problems a sequence of $n$ independent buyers arrive one-by-one. However, in SPP we want to maximize the revenue and in PI we want to maximize the welfare. Interestingly, for known buyer value distributions, both these problems are equivalent [CFPV19], but this reduction requires virtual-value distributions and doesn’t work for unknown distributions. In the bandit model, PI and SPP behave very differently. For instance, [GKSW24] recently obtained $\widetilde{O}(\sqrt{T})$ regret algorithm for PI with general distributions, whereas an $\Omega(T^{2/3})$ regret lower bound exists for SPP [LSTW23b]. Our $\widetilde{O}(\sqrt{T})$ regret results crucially rely on half-concavity of regular distributions.
Regular Distributions and Learning. Myerson’s regularity of distributions has been greatly studied since it was introduced in [Mye81]. In particular, it is a standard assumption in learning theory for auctions [DRY15, CR14, HMR18, DHP16, RS16, GHZ19]. For basic properties of regular distribution, and its important subclass of Monotone Hazard Rate distributions, see the books [Har13, Rou17]. A recent paper of [LSTW23b], which studies pricing query complexity of the single-buyer single-item problem, proves an interesting “relative flatness” property of regular distributions in the value space. Roughly it says that if the revenue curve has nearly the same value at 4 different equidistant points then there cannot be a high revenue point in between. Comparing it to our idea of half-concavity, the two properties are in general incomparable as no one implies the other. However, we found our half-concavity to be more intuitive and convenient to work with, since proofs based on relative flatness often lead to a long case analysis.
2 A Single Buyer with Regular Distribution
In this section, we present an $\widetilde{O}(\sqrt{T})$ regret algorithm for BSPP in the special case of a single buyer, where we call the problem Bandit Posted Pricing. Moreover, we focus on the case when the buyer’s value distribution is regular, which is a standard assumption in economics [Mye81].
Definition 2.1 (Regularity).
Distribution $\mathcal{D}$ with CDF $F(x)$ and PDF $f(x)$ is called regular when $\phi(v)~{}:=~{}v-\frac{1-F(v)}{f(v)}$ is monotone non-decreasing, or equivalently, its revenue curve $R_{q}(q)$ is concave in the quantile space, where
$R_{q}(q)~{}:=~{}q\cdot F^{-1}(1-q).$
In the Bandit Posted Pricing problem there is a single regular buyer with an unknown regular value distribution $\mathcal{D}$ having support $[0,1]$ and CDF $F(x)$. Our goal is to approach the optimal price $p^{*}$ that maximizes the revenue function $R(p):=p\cdot(1-F(p))$ in the the following bandit learning game over $T$ days: On day $t\in\{1,\ldots,T\}$, we post a price $p_{t}\in[0,1]$ and the environment draws $v_{t}\sim\mathcal{D}$. Our reward is $p_{t}\cdot\textbf{1}_{v_{t}\geq p_{t}}$, and the goal is to minimize in expectation the total regret: $T\cdot R(p^{*})-\sum_{t\in[T]}p_{t}\cdot\textbf{1}_{v_{t}\geq p_{t}}.$ Our $\widetilde{O}(\sqrt{T})$ regret algorithm for this problem uses “half-concavity”.
2.1 Half-Concavity
Our $\widetilde{O}(\sqrt{T})$ regret algorithm works beyond regular distributions, for the class of half-concave distributions. We first give the definition of half-concavity:
Definition 2.2 (Half-Concavity).
A function $R(x):\mathbb{R}\rightarrow\mathbb{R}$ is half-concave in interval $[\ell,r]\subseteq[0,1]$ if the following conditions hold:
  1. (i)
    $R(x)$ is single-peaked in $[\ell,r]$, i.e., $\exists p^{*}\in[\ell,r]$ satisfying that $R(x)$ is non-decreasing in $[\ell,p^{*}]$ and non-increasing in $[p^{*},r]$.
  2. (ii)
    $R(x)$ is 1-Lipschitz in $[\ell,p^{*}]$.
  3. (iii)
    $R(x)$ is concave in $[\ell,p^{*}]$.
Figure 1: An example of a half-concave function, where the function is Lipschitz and concave before $p^{*}$.
The concept of half-concavity implies that a function has nice properties properties on one side of the maximum (see Figure 1), rendering it learnable from this advantageous side. In the definition of half-concavity, both Lipschitzness and some kind of concavity are vital to ensure learnability. Without Lipschitzness, a function becomes unlearnable when $p^{*}$ is exceedingly close to 00 due to the inability of accurately detecting the value of $p^{*}$. On the other hand, without concavity, simple examples show that a regret of $\Omega(T^{2/3})$ is inevitable. For instance, a revenue function characterized by multiple peaks where we can only ascertain the value of $p^{*}$ when we examine each peak with a sufficient number of samples. We note that related works, such as [CP14, MCP14], address bandit problems with assumptions of unimodularity or Lipschitzness. However, these works only provide instance dependent regret bounds, which are at least $\omega(\sqrt{T})$ in the worst case, as they lack at least one of the key assumptions of Lipschitzness and concavity.
Given both Lipschitzness and (half) concavity, the following theorem gives $\widetilde{O}(\sqrt{T})$ regret.
Theorem 2.3.
For Bandit Posted Pricing with a single buyer, if the revenue function $R(p):=p\cdot(1-F(p))$ is half-concave, then there exists an algorithm with $O(\sqrt{T}\log T)$ regret.
The main application of 2.3 is for regular distributions due to the following lemma.
Lemma 2.4.
Let $F(x)$ be the CDF of a regular distribution with support in $[0,1]$, and $R(p):=p\cdot(1-F(p))$ be the expected revenue on playing price $p$. Then, $R(p)$ is half-concave in $[0,1]$.
Before proving this lemma, we observe the following immediate corollary.
Corollary 2.5 (Corollary of 2.3 and 2.4).
For Bandit Posted Pricing with a single buyer having regular valuation distribution, there exists an algorithm with $O(\sqrt{T}\log T)$ regret.
Proof of 2.4.
We first provide a proof assuming the PDF $f(p)$ is non-zero and differentiable. We will prove the three properties of half-concavity one-by-one.
(i) Single-peakedness: By definition, revenue function of a regular distribution is concave in the quantile space. Hence, it’s single-peaked in the quantile space. Suppose $q^{*}$ is the quantile that achieves the maximum, then choosing price $p^{*}=F^{-1}(1-q^{*})$ proves single-peakedness of $R(p)$.
(ii) Lipschitzness: For $0\leq a\leq b\leq p^{*}$, we have
$0~{}\leq~{}R(b)-R(a)~{}=~{}(b-a)-F(b)\cdot b+F(a)\cdot a~{}\leq~{}b-a,$
where the first inequality follows from single-peakedness and that $a\leq b\leq p^{*}$, and the second inequality uses $F(b)\geq F(a)\geq 0$ and that $b\geq a\geq 0$.
(iii) Half-concavity: We first provide a direct proof via the monotonicity of virtual valuation $\phi$.
Recall that $\phi(p):=p-\frac{1-F(p)}{f(p)}$ is non-decreasing for regular distribution, so we have
$\phi^{\prime}(p)~{}=~{}1-\frac{-f^{2}(p)-(1-F(p))\cdot f^{\prime}(p)}{f^{2}(p)}~{}\geq~{}0.$
Rearranging the above inequality,
$0\displaystyle 2f^{2}(p)+f^{\prime}(p)\cdot(1-F(p))~{}\geq~{}$ 0 . (1)
Consider the derivative and the second order derivative of $R(p)=p\cdot(1-F(p))$, i.e.,
$R^{\prime}(p)~{}=~{}1-F(p)-p\cdot f(p)\qquad\text{and}\qquad R^{\prime\prime}(p)~{}=~{}-2f(p)-p\cdot f^{\prime}(p).$
For half-concavity, we prove that $R^{\prime\prime}(p)\leq 0$ for every $p\in[0,p^{*}]$. Consider two cases:
  • Case 1: $f^{\prime}(p)>0$. In this case, $R^{\prime\prime}(p)=-2f(p)-p\cdot f^{\prime}(p)\leq 0$ immedately holds.
  • Case 2: $f^{\prime}(p)\leq 0$. In this case, we have
    $R^{\prime\prime}(p)~{}$ $=~{}-2f(p)-p\cdot f^{\prime}(p)$
    $=~{}\frac{1}{f(p)}\cdot\left(-2f^{2}(p)-f(p)\cdot p\cdot f^{\prime}(p)\right)$
    $\leq~{}\frac{1}{f(p)}\cdot\left(f^{\prime}(p)\cdot(1-F(p))-f(p)\cdot p\cdot f^{\prime}(p)\right)$
    $=~{}\frac{f^{\prime}(p)}{f(p)}\cdot\big{(}1-F(p)-p\cdot f(p)\big{)}~{}\leq~{}0,$
    where the first inequality is from (1) and the second inequality uses $f^{\prime}(p)\leq 0$ along with the fact that $R^{\prime}(p)\geq 0$ when $p\in[0,p^{*}]$.
The above proof requires the PDF $f(p)$ to be non-zero and differentiable. Below we provide an alternate proof via the concavity in the quantile space, which bypasses this assumption for $f(p)$. We keep both the proofs here since the reader may find one proof more instructive than the other.
An alternate proof for half-concavity. We prove half-concavity by contradiction. Assume there exist $0\leq a<b<c\leq p^{*}$ satisfying $c=b+t(b-a)$ and $R(c)>R(b)+t\cdot\big{(}R(b)-R(a)\big{)}$.
Define $q(x):=1-F(x)$ and $\overline{q}:=q(b)+t\cdot\big{(}q(b)-q(a)\big{)}$. We first show that $\overline{q}>q(c)$: Regularity implies that the revenue function is concave in the quantile space. Hence, for all $q^{\prime}\geq\overline{q}$,
$R(q^{-1}(q^{\prime}))~{}\leq~{}R(b)+\frac{q(b)-q^{\prime}}{q(a)-q(b)}\cdot(R(b)-R(a))~{}\leq~{}R(b)+t\cdot(R(b)-R(a))~{}<~{}R(c).$
For every $q^{\prime}\geq\bar{q}$, we have $q^{\prime}\neq q(c)$. So there must be $q(c)<\bar{q}$. This gives the desired contradiction:
$R(c)$ $~{}>~{}R(b)+t\cdot(R(b)-R(a))$
$~{}=~{}b\cdot q(b)+t\cdot\big{(}b\cdot q(b)-a\cdot q(a)\big{)}$
$~{}\geq~{}b\cdot q(b)+t\cdot\big{(}b\cdot q(b)-a\cdot q(a)\big{)}-t(t+1)(b-a)(q(a)-q(b))$
$~{}=~{}\big{(}(t+1)b-ta\big{)}\cdot\big{(}(t+1)q(b)-tq(a)\big{)}$
$~{}=~{}c\cdot\overline{q}~{}\geq~{}c\cdot q(c)~{}=~{}R(c).\qed$
2.2 $\widetilde{O}(\sqrt{T})$ Regret for Half-Concave Functions: Proof of 2.3
Our algorithm relies on a subroutine that takes a parameter $\epsilon$ and a confidence interval $[\ell,r]\ni p^{*}$ as input, and then after $\widetilde{O}(\epsilon^{-2})$ rounds it generates with high probability a new confidence interval $[\ell^{\prime},r^{\prime}]$ that contains the optimal price $p^{*}$ and every price in $[\ell^{\prime},r^{\prime}]$ has $1$-day regret bounded by $\epsilon$. Our final algorithm runs in $O(\log T)$ phases, where in each phase we call the sub-routine with the parameter $\epsilon$, and $\epsilon$ is halved in the next phase. The sub-routine is captured by the following lemma, which is the heart of the proof and will be proved in Section 2.3 using half-concavity of $R(p)$.
Lemma 2.6.
Let $R(p)$ be a half-concave revenue function defined in $[0,1]$. Given $\epsilon\geq{1}/{\sqrt{T}}$ and $[\ell,r]\ni p^{*}$, there exists an algorithm that tests $O\big{(}\epsilon^{-2}{\log^{2}T}\big{)}$ rounds with prices inside $[\ell,r-\frac{1}{T^{100}}]$ and outputs an interval $[\ell^{\prime},r^{\prime}]\subseteq[\ell,r]$ satisfying with probability $1-T^{-5}$ that
  • $p^{*}\in[\ell^{\prime},r^{\prime}]$
  • $R(x)\geq R(p^{*})-\epsilon$ for any $x\in[\ell^{\prime},r^{\prime}-\frac{1}{T^{100}}]$.
Input: Hidden Revenue Function $R(p)$, time horizon $T$
1 Let $\epsilon_{1}\leftarrow 1,[\ell_{1},r_{1}]\leftarrow[0,1]$, and $i\leftarrow 1$.
2 while $\epsilon_{i}>\frac{\log T}{\sqrt{T}}$ do
3       Run the algorithm described in 2.6 with $\epsilon_{i},[\ell_{i},r_{i}]$ as input, and get $[\ell^{\prime},r^{\prime}]$.
4       Let $[\ell_{i+1},r_{i+1}]\leftarrow[\ell^{\prime},r^{\prime}]$ and $\epsilon_{i+1}\leftarrow\frac{1}{2}\epsilon_{i}$.
5       Let $i\leftarrow i+1$.
Finish remaining rounds with any price in $[\ell_{i},r_{i}-{T^{-100}}]$.
ALGORITHM 1 $O(\sqrt{T}\log T)$ Regret Algorithm
Given 2.6, our algorithm for 2.3 is natural and has a simple proof.
Proof of 2.3.
We claim that with probability at least $1-T^{-4}$, the regret of Algorithm 1 is $O(\sqrt{T}\log T)$. Algorithm 1 uses 2.6 for multiple times with a halving error parameter. Assume Algorithm 1 ends with $i=k+1$, i.e., the while loop runs $k$ times. Since we obtain $\epsilon_{i+1}=\epsilon_{i}/2$ and $\epsilon_{k}>\frac{\log T}{\sqrt{T}}$ holds, there must be $k=O(\log T)$.
For $i\in[k]$, let $\mathsf{Alg}_{i}$ represent the corresponding algorithm we call when using 2.6 with $\epsilon_{i},[\ell_{i},r_{i}]$ as the input. To use 2.6 we need to verify that $p^{*}\in[\ell_{i},r_{i}]$ for all $i\in[k]$. The condition $p^{*}\in[\ell_{1},r_{1}]=[0,1]$ clearly holds. For $i=2,3,\cdots,k$, the condition $p^{*}\in[\ell_{i},r_{i}]$ is guaranteed by 2.6 when calling $\mathsf{Alg}_{i-1}$. The failing probability of 2.6 is $T^{-5}$. By the union bound, with probability $1-k\cdot T^{-5}>1-T^{-4}$, 2.6 always holds in Algorithm 1.
Now we prove the regret of Algorithm 1. 2.6 guarantees that $R(p^{*})-R(p)\leq\epsilon_{i}$ holds for all $p\in[\ell_{i+1},r_{i+1}-{T^{-100}}]$, while when calling $\mathsf{Alg}_{i+1}$, only the prices in $[\ell_{i+1},r_{i+1}-{T^{-100}}]$ is tested. Therefore, the total regret of Algorithm 1 can be bounded by
$1\cdot O\big{(}\epsilon^{-2}_{1}\log^{2}T\big{)}~{}+~{}\sum_{i=2}^{k}\epsilon_{i-1}\cdot O(\epsilon^{-2}_{i}\log^{2}T)~{}+~{}T\cdot\epsilon_{k}~{}=~{}O(\sqrt{T}\log T).$
In this expression, the first term is the regret of $\mathsf{Alg}_{1}$. The second term is the sum of the regret from $\mathsf{Alg}_{2},\cdots,\mathsf{Alg}_{k}$. The third term is the regret for the remaining rounds.
Finally, we also need to check that Algorithm 1 uses no more than $T$ rounds. 2.6 suggests that $\mathsf{Alg}_{i}$ runs $O({\epsilon^{-2}_{i}}{\log^{2}T})$ rounds. So the total number of rounds in the while loop is $\sum_{i\in[k]}O({\epsilon^{-2}_{i}}{\log^{2}T})~{}=~{}O(T).$ Therefore, Algorithm 1 is feasible. ∎
2.3 Main Sub-Routine: Proof of 2.6 via Half-Concavity
In this section, we present the main sub-routine of our bandit algorithm. It utilizes half-concavity to generate a confidence interval for the optimal price $p^{*}$ in $\widetilde{O}(\frac{1}{\epsilon^{2}})$ rounds while ensuring that every price in the interval is $\epsilon$-optimal.
Algorithm Overview.
The algorithm contains two major steps (also see Figure 2):
  • Step 1: Find an approximation $\hat{p}$ such that $R(p^{*})-R(\hat{p})\leq\frac{\epsilon}{2}$ via half-concavity. This step gives a sufficiently precise price estimate approximating $p^{*}$.
  • Step 2: Given $\hat{p}$, construct a new confidence interval $[\ell^{\prime},r^{\prime}]$ via binary search. As $R(p)$ exhibits a single-peak property, we can use $R(\hat{p})$ as a benchmark and implement a standard binary search algorithm to identify the leftmost point $\ell^{\prime}$ and the rightmost point that are $\frac{\epsilon}{2}$-close to $R(\hat{p})$. These two endpoints describe the desired new confidence interval.
We note an important detail in Step 2: the Lipschitzness assumption in the range $[p^{*},r]$ is absent. Consequently, the binary search algorithm fails to provide a satisfactory loss guarantee for a small tail of the new confidence interval. This introduces the non-standard error factor ${T^{-100}}$ as discussed in 2.6.
Figure 2: The two main steps of 2.6
2.3.1 Step 1: Find $\hat{p}$ to Approximate $p^{*}$
The first step of the sub-routine is to find $\hat{p}$ that approximates $p^{*}$. Specifically, we want to prove the following lemma:
Lemma 2.7.
Assume $R(p)$ is a half-concave revenue function defined in $[0,1]$. Given interval $[\ell,r]$ and error parameter $\epsilon>{1}/{\sqrt{T}}$, there exists an algorithm that tests $O({\epsilon^{-2}}{\log^{2}T})$ rounds with prices inside $[\ell,r-{T^{-100}}]$ and outputs $\hat{p}\in[\ell,r-{T^{-100}}]$ satisfying with probability at least $1-T^{-6}$ that $\max_{p\in[\ell,r]}R(p)-R(\hat{p})<\frac{\epsilon}{2}$.
When we have the assumption $p^{*}\in[\ell,r]$ from 2.6, 2.7 implies we are able to find $\hat{p}$ such that $R(p^{*})-R(\hat{p})\leq\frac{\epsilon}{2}$ in the first step. We first give the pseudo-code of the algorithm required by 2.7 and its high-level idea.
Input: Hidden Revenue Function $R(p)$, Interval $[\ell,r]$, Error Parameter $\epsilon$
1 Let $\delta=\frac{\epsilon}{100}$ be the scaled error parameter, $\ell_{1}\leftarrow\ell,r_{1}\leftarrow r$, $i=1$.
2 while $r_{i}-\ell_{i}>\delta$ do
3       Let $a_{i}\leftarrow({2\ell_{i}+r_{i}})/{3}$ and $b_{i}\leftarrow({\ell_{i}+2r_{i}})/{3}$.
4       Test $C\cdot{\delta^{-2}}{\log T}$ rounds with price $p=a_{i}$. Let $\hat{R}(a_{i})$ be the average reward.
5       Test $C\cdot{\delta^{-2}}{\log T}$ rounds with price $p=b_{i}$. Let $\hat{R}(b_{i})$ be the average reward.
6       if $\hat{R}(a_{i})<\hat{R}(b_{i})-2\delta$ then
7            Let $\ell_{i+1}\leftarrow a_{i},r_{i+1}\leftarrow r_{i}$
8      else
9            Let $\ell_{i+1}\leftarrow\ell_{i},r_{i+1}\leftarrow b_{i}$
10      $i\leftarrow i+1$
11Test $C\cdot{\delta^{-2}}{\log T}$ rounds with price $p=\ell_{i}$. Let $\hat{R}(\ell_{i})$ be the average reward.
12 Let $\hat{p}=\arg\max_{p\in P}\hat{R}(p)$, where $P=\{p:p\text{ is tested}\}$.
Output: $\hat{p}$.
ALGORITHM 2 Finding $\hat{p}$
Algorithm 2 is a recursive algorithm that runs $O(\log\frac{1}{\delta})$ rounds. In each round, the algorithm tests the two points $a,b$ that divide the current interval $[\ell,r]$ into thirds. Each price is tested for $\widetilde{O}(\frac{1}{\delta^{2}})$ rounds, where $\delta={\epsilon}/{100}$ is the scaled error parameter. A standard concentration inequality guarantee both $|R(a)-\hat{R}(a)|\leq\delta$ and $|R(b)-\hat{R}(b)|\leq\delta$ hold. Then, the algorithm drops one third of the interval according to the test results. There are two different cases (see Figure 3):
Figure 3: Two cases of Step 1 as discussed in Algorithm 2.
Case 1: $\hat{R}(a)<\hat{R}(b)-2\delta$. The inequality is sufficient to show that $R(a)<R(b)$. Since function $R(p)$ is single-peaked, there must be $p^{*}\geq a$ and the sub-interval $[\ell,a]$ can be dropped.
Case 2: $\hat{R}(a)\geq\hat{R}(b)-2\delta$. This inequality implies $R(b)-R(a)$ is sufficiently small. In this case, observe that if $p^{*}\in[b,r]$, the value of $R(p^{*})$ can’t be much better than $R(b)$, because the concavity in $[a,p^{*}]$ guarantees that $R(p^{*})-R(b)\leq R(b)-R(a)$. Therefore, the sub-interval $[b,r]$ can be dropped because we don’t need to further test it.
Finally, the recursive algorithm stops when the length of the remaining interval is bounded by $\delta$, and the Lipschitzness guarantees the final interval is good.
Now, we give the formal proof of 2.7.
Proof of 2.7.
We first show that Algorithm 2 tests $O(\frac{\log^{2}T}{\epsilon^{2}})$ prices in $[\ell,r-{T^{-100}}]$.
The main body of Algorithm 2 is a while loop that maintains an interval to be tested. Assume the while loop stops when $i=k+1$. Then, for $i\in[k]$, the length of $[\ell_{i+1},r_{i+1}]$ is two-thirds of $[\ell_{i},r_{i}]$, and we have $r_{k}-\ell_{k}>\delta$. Therefore, $k=O(\log\frac{1}{\delta})=O(\log T)$. For interval $[\ell_{i},r_{i}]$, two prices $a_{i},b_{i}$ are tested for $O({\delta^{-2}}{\log T})$ times, so in total Algorithm 2 tests $O({\delta^{-2}}{\log^{2}T})$ rounds. To show every tested price is in $[\ell,r-{T^{-100}}]$, let $p$ be a price tested in Algorithm 2. $p\in[\ell,r]$ directly follows from the algorithm. For the inequality $p\leq r-{T^{-100}}$, observe that the largest tested price must be $p=b_{i}$ for some $i\in[k]$, and $b_{i}\leq r-{T^{-100}}$ is guaranteed by the fact that $r_{i}-b_{i}\geq\frac{1}{3}\delta\gg{T^{-100}}$.
It only remains to show $\max_{p\in[\ell,r]}-R(\hat{p})\leq\frac{\epsilon}{2}$. Define $p^{*}_{\ell,r}:=\arg\max_{p\in[\ell,r]}R(p)$ to be the optimal price in $[\ell,r]$. We will show that $R(p^{*}_{\ell,r})-R(\hat{p})\leq 6\delta<\frac{\epsilon}{2}$ holds with probability $1-T^{-6}$. We first need the following claim, which follows from standard concentration inequalities:
Claim 2.8.
In Algorithm 2, $|\hat{R}(p)-R(p)|\leq\delta$ simultaneously holds for every tested price $p$ with probability at least $1-T^{-6}$.
Proof.
For a single tested price $p$, $\hat{R}(p)$ is estimated by calculating the average of $N=C\cdot\frac{\log T}{\delta^{2}}$ samples. By Hoeffding’s Inequality,
$\mbox{\rm\bf Pr}\left[|\hat{R}(p)-R(p)|>\delta\right]~{}\leq 2\exp\left(-2N\delta^{2}\right)~{}=~{}2T^{-2C}~{}<~{}T^{-7}.$
The last inequality holds when $C$ is a constant greater than $4$. Then, we have $|\hat{R}(p)-R(p)|\leq\delta$ holds with probability $1-T^{-7}$ for a single tested price $p$.
Notice that Algorithm 2 can’t test more than $T$ prices. By the union bound, $|\hat{R}(p)-R(p)|\leq\delta$ simultaneously holds for all tested prices with probability $1-T^{-6}$. ∎
Next, we use 2.8 to show $R(p^{*}_{\ell,r})-R(\hat{p})\leq 6\delta$. We consider the following two cases:
The first case is that there exists $i\in[k]$, such that $p^{*}_{\ell,r}$ falls in $[\ell_{i},r_{i}]$ but not in $[\ell_{i+1},r_{i+1}]$. In this case, there must be $\hat{R}(a_{i})\geq\hat{R}(b_{i})-2\delta$ and $[b_{i},r_{i}]$ is dropped. If not, we have $\hat{R}(a_{i})\geq\hat{R}(b_{i})-2\delta$ and $[\ell_{i},a_{i}]$ is dropped. However, the inequality $\hat{R}(a_{i})\geq\hat{R}(b_{i})-2\delta$ together with 2.8 gives $R(a_{i})<R(b_{i})$. Then, the single-peakedness of $R(p)$ gives $p^{*}_{\ell,r}\geq a_{i}$, which is in contrast to the assumption that $p^{*}_{\ell,r}\notin[\ell_{i+1},r_{i+1}]$. Therefore, we have $\hat{R}(a_{i})\geq\hat{R}(b_{i})-2\delta$, and the assumption $p^{*}_{\ell,r}\notin[\ell_{i+1},r_{i+1}]$ implies $p^{*}_{\ell,r}\in[b_{i},r_{i}]$. Then, the concavity of $R(p)$ in $[a_{i},p^{*}_{\ell,r}]$ guarantees
$R(p^{*}_{\ell,r})-R(b_{i})~{}\leq~{}R(b_{i})-R(a_{i})~{}\leq~{}(\hat{R}(b_{i})+\delta)-(\hat{R}(a_{i})-\delta)~{}\leq~{}4\delta.$
The second case is that the desired $i$ in the first case doesn’t exist. In this case, the only possibility is that $p^{*}_{\ell,r}\in[\ell_{k+1},r_{k+1}]$. Then, the 1-Lipschitzness of function $R(p)$ guarantees that $R(p^{*}_{\ell,r})-R(\ell_{k})\leq\delta$.
In both cases, we find a tested price $p^{\prime}$ satisfying $R(p^{\prime})\geq R(p^{*}_{\ell,r})-4\delta$. Then,
$R(\hat{p})~{}\geq~{}\hat{R}(\hat{p})-\delta~{}\geq\hat{R}(p^{\prime})-\delta~{}\geq~{}R(p^{\prime})-2\delta~{}\geq R(p^{*}_{\ell,r})-6\delta.\qed$
2.3.2 Step 2: Generating New Confidence Interval
Upon executing Algorithm 2, we obtain an approximate optimal price $\hat{p}$. The subsequent stage of the algorithm uses $R(\hat{p})$ as a benchmark to establish the new upper and lower bounds of the confidence interval. Given that the function $R(p)$ is single-peaked, two independent binary search algorithms suffice to independently determine these new upper and lower bounds, leading to the following 2.9:
Lemma 2.9.
Assume $R(p)$ is a half-concave revenue function defined in $[0,1]$. Given interval $[\ell,r]\ni p^{*}$, error parameter $\epsilon>\frac{1}{\sqrt{T}}$, near-optimal price $\hat{p}\in[\ell,r-{T^{-100}}]$ that satisfies $R(p^{*})-R(\hat{p})\leq\frac{\epsilon}{2}$, there exists an algorithm that tests $O(\epsilon^{-2}\log^{2}T)$ rounds with prices in $[\ell,r-{T^{-100}}]$ and outputs $[\ell^{\prime},r^{\prime}]\subseteq[\ell,r]$ satisfying with probability $1-T^{-6}$ that:
  • $p^{*}\in[\ell^{\prime},r^{\prime}]$.
  • For $p\in[\ell^{\prime},r^{\prime}-{T^{-100}}]$, we have $R(p^{*})-R(p)\leq\epsilon$.
Combining 2.7 and 2.9 proves 2.6:
Proof of 2.6.
We first run Algorithm 2 to get $\hat{p}$, and then run the algorithm described in 2.9 with $\hat{p}$ being part of the input. Combining 2.7 and 2.9 with union bound guarantees that we have the desired output with success probability at least $1-2T^{-6}>1-T^{-5}$. ∎
It only remains to prove 2.9. We first give the following Algorithm 3, and then show Algorithm 3 is the desired algorithm for 2.9.
Input: Hidden Revenue Function $R(p)$, Interval $[\ell,r]$, Error Parameter $\epsilon$, Near-Optimal price $\hat{p}$
1 Let $\delta=\frac{\epsilon}{100}$ be the scaled error parameter.
// First binary search to determine $\ell^{\prime}$
2 Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p=\hat{p}$. Let $\hat{R}(\hat{p})$ be the average reward.
3 Let $\ell_{b1}\leftarrow\ell,r_{b1}\leftarrow\hat{p}$.
4 while $r_{b1}-\ell_{b1}\geq{T^{-100}}$ do
5       Let $m=\frac{\ell_{b1}+r_{b1}}{2}$.
6       Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p=m$. Let $\hat{R}(m)$ be the average reward.
7       if $\hat{R}(m)<\hat{R}(\hat{p})-2\delta$ then Update $r_{b1}\leftarrow m$ else Update $\ell_{b1}\leftarrow m$;
8
9Let $\ell^{\prime}\leftarrow\ell_{b1}$
// Second binary search to determine $r^{\prime}$
10 Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p=r-{T^{-100}}$. Let $\hat{R}(r-{T^{-100}})$ be the average reward.
11 if $\hat{R}(r-{T^{-100}})\geq\hat{R}(\hat{p})-2\delta$ then
       // Special Case: original upper bound is good
12       Let $r^{\prime}\leftarrow r$
13else
14       Let $\ell_{b2}\leftarrow\hat{p},r_{b2}\leftarrow r-{T^{-100}}$.
15       while $r_{b2}-\ell_{b2}\geq{T^{-100}}$ do
16             Let $m=\frac{\ell_{b2}+r_{b2}}{2}$.
17             Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p=m$. Let $\hat{R}(m)$ be the average reward.
18             if $\hat{R}(m)<\hat{R}(\hat{p})-2\delta$ then Update $r_{b2}\leftarrow m$ else Update $\ell_{b2}\leftarrow m$;
19
20      Let $r^{\prime}\leftarrow r_{b2}$.
Output: $[\ell^{\prime},r^{\prime}]$.
ALGORITHM 3 Getting $[\ell^{\prime},r^{\prime}]$ via Binary Search
We first give the following concentration bound for Algorithm 3:
We first give the following claim:
Claim 2.10.
In Algorithm 3, $|\hat{R}(p)-R(p)|\leq\delta|$ simultaneously holds for every tested price $p$ with probability at least $1-T^{-6}$.
Proof.
For a single tested price $p$, $\hat{R}(p)$ is estimated by calculating the average of $N=C\cdot\frac{\log T}{\delta^{2}}$ samples. By Hoeffding’s Inequality,
$\mbox{\rm\bf Pr}\left[|\hat{R}(p)-R(p)|>\delta\right]~{}\leq~{}2\exp\left(-2N\delta^{2}\right)~{}=~{}2T^{-2C}~{}<~{}T^{-7}.$
The last inequality holds when $C$ is a constant greater than $4$. Then, we have $|\hat{R}(p)-R(p)|\leq\delta$ holds with probability $1-T^{-7}$ for a single tested price $p$.
Notice that Algorithm 3 can’t test more than $T$ prices. By the union bound, $|\hat{R}(p)-R(p)|\leq\delta$ simultaneously holds for all tested prices with probability $1-T^{-6}$. ∎
Now we prove 2.9 assuming 2.10 holds.
Proof of 2.9.
We show Algorithm 3 satisfies the statements in 2.9. Our proof starts from verifying Algorithm 3 runs $O(\frac{\log T}{\delta^{2}})$ rounds with prices in $[\ell,r-{T^{-100}}]$. For the number of tested rounds, Algorithm 3 runs two standard binary searches. In one binary search, the subroutine tests one price with $C\cdot\frac{\log T}{\delta^{2}}$ rounds, and the binary search stops when the length of the interval is less than ${T^{-100}}$. Therefore, The total number of rounds is $C\cdot\frac{\log T}{\delta^{2}}\cdot O(\log T^{100})=O(\frac{\log^{2}T}{\epsilon^{2}})$. For the range constraint of all tested prices, Algorithm 3 directly guarantees that every tested price falls in $[\ell,r-{T^{-100}}]$.
Next, we show $p^{*}\in[\ell^{\prime},r^{\prime}]$. We first prove $p^{*}\geq\ell^{\prime}$. In Algorithm 3, the only way we update $\ell^{\prime}$ is that we observe $\hat{R}(m)<\hat{R}(\hat{p})-2\delta$ in the first binary search, and $\ell^{\prime}$ is updated to at least $m$. This is feasible because 2.10 gives $R(m)<R(\hat{p})\leq R(p^{*})$, and the single-peakedness of $R(p)$ guarantees $m\leq p^{*}$. The proof of $p^{*}\leq r^{\prime}$ is symmetric.
Finally, we show $R(p)\geq R(p^{*})-\epsilon$ for all $p\in[\ell^{\prime},r^{\prime}-{T^{-100}}]$. Function $R(p)$ is single-peaked, so it is sufficient to prove that $R(\hat{p})-R(\ell^{\prime})\leq 5\delta<\frac{\epsilon}{2}$ and $R(\hat{p})-R(r^{\prime}-{T^{-100}})\leq 5\delta<\frac{\epsilon}{2}$. Then, combining these two inequalities with the assumption $R(p^{*})-R(\hat{p})\leq\frac{\epsilon}{2}$ gives the desired statement.
We first prove that $R(\hat{p})-R(\ell^{\prime})\leq 5\delta$. The binary search subroutine guarantees that $\hat{R}(r_{b1})\geq\hat{R}(\hat{p})-2\delta$ after the binary search loop is finished. Then, 2.10 ensures that $R(\hat{p})-R(r_{b1})\leq 4\delta$. Since $\ell^{\prime}$ is set to be $\ell_{b1}>r_{b1}-{T^{-100}}$, the 1-Lipschitzness of $R(p)$ in $[\ell_{b1},\hat{p}]$ guarantees that
$R(\hat{p})-R(\ell^{\prime})~{}=~{}R(\hat{p})-R(r_{b1})+R(r_{b1})-R(\ell_{b1})~{}\leq~{}4\delta+{T^{-100}}~{}\leq~{}5\delta.$
For the inequality $R(\hat{p})-R(r^{\prime}-{T^{-100}})\leq 5\delta$, a symmetric proof for the second binary search subroutine gives $R(\hat{p})-R(\ell_{b2})\leq 4\delta$, while $r^{\prime}$ is set to be $r_{b2}<\ell_{b2}+{T^{-100}}$. If $p^{*}<r^{\prime}-{T^{-100}}$, the single-peakedness of $R(p)$ gives
$R(r^{\prime}-{T^{-100}})~{}\geq~{}R(\ell_{b2})~{}\geq~{}R(\hat{p})-4\delta.$
Otherwise, the 1-Lipschitzness gives
$R(r^{\prime}-{T^{-100}})~{}\geq~{}R(\ell_{b2})-(\ell_{b2}-r_{b2}+{T^{-100}})~{}\geq~{}R(\hat{p})-4\delta-{T^{-100}}>R(\hat{p})-5\delta.\qed$
3 $n$ Buyers with Regular Distributions
In this section, we provide the $\widetilde{O}(\mathsf{poly}(n)\sqrt{T})$ regret algorithm for BSPP with $n$ buyers having regular distributions. In this problem we have $n$ buyers with independent regular value distributions $\mathcal{D}_{1},\cdots,\mathcal{D}_{n}$. The seller posts prices $p_{1},\cdots,p_{n}$, and the first buyer with value $v_{i}\sim\mathcal{D}_{i}\geq p_{i}$ gets the item and pays $p_{i}$. Our goal is to approach the optimal prices $p^{*}_{1},\cdots,p^{*}_{n}$ that maximizes the expected revenue $R(p_{1},\cdots,p_{n}):=p_{i}\cdot\mbox{\rm\bf Pr}\left[i\text{ gets the item}\right]$ via a bandit learning game over $T$ days: On day $t\in[T]$ we post a price vector $(p^{(t)}_{1},\cdots,p^{(t)}_{n})$ and the environment draws $v^{(t)}_{1}\sim\mathcal{D}_{1},\cdots,v^{(t)}_{n}\sim\mathcal{D}_{n}$. We only observe reward $\text{Rev}_{t}=\sum_{i\in[n]}p^{(t)}_{i}\cdot\mathbf{1}[p^{(t)}_{i}\geq v^{(t)}_{i}\land\forall j<i,p^{(t)}_{j}<v^{(t)}_{j}]$. Our goal is to minimize the total regret $T\cdot R(p^{*}_{1},\cdots,p^{*}_{n})-\sum_{t\in[T]}\text{Rev}_{t}$ in expectation. Our main result is:
Theorem 3.1.
For BSPP with $n$ buyers having regular distributions, there exists an algorithm with $O(n^{2.5}\sqrt{T}\log T)$ regret.
3.1 Proof Overview
In this section, we provide an overview of our algorithm for sequential buyers. The main structure of our algorithm follows the structure of the single-buyer algorithm. We maintain $n$ confidence intervals $[\ell_{1},r_{1}],\cdots,[\ell_{n},r_{n}]$ for the optimal prices $p^{*}_{1},\cdots,p^{*}_{n}$. The core of the algorithm is to design a sub-routine that uses $\widetilde{O}({\mathsf{poly}(n)}/{\epsilon^{2}})$ rounds to update the confidence intervals to $[\ell^{\prime}_{1},r^{\prime}_{1}],\cdots,[\ell^{\prime}_{n},r^{\prime}_{n}]$, such that playing any prices inside new confidence intervals has regret at most $\epsilon$. Then, calling this sub-routine $O(\log T)$ times with halving error parameter $\epsilon$ suffices to obtain $\widetilde{O}(\mathsf{poly}(n)\sqrt{T})$ regret.
The core sub-routine contains two steps. The first step is to get a group of near-optimal prices $\hat{p}_{1},\cdots,\hat{p}_{n}$. The second step is to use prices $\hat{p}_{1},\cdots,\hat{p}_{n}$ as a benchmark to find new upper- and lower- confidence bounds.
Step 1: Finding $\hat{p}_{i}$ with a Different Revenue Function. Comparing to the single-buyer setting, the revenue function for sequential buyers is different. In particular, when we want to determine $\hat{p}_{i}$, the revenue function for buyer $i$ looks like
$R(p)=p\cdot(1-F(p))+C\cdot F(p),$
where the constant $C$ represents the expected revenue from the buyers $i+1,\cdots,n$. Our goal is to find a near-optimal price $\hat{p}_{i}$ for this new revenue function $R(p)$.
The major challenge in this step is that function $R(p)$ is not always half-concave: it is half-concave in $[C,1]$, while in $[0,C]$ it is increasing but not necessary concave. This prevents us from directly applying the single-buyer algorithm. To fix this, an idea is to estimate the value of $C$ to determine the half-concave region. We show that $C$ can be estimated accurately enough in some of the cases, where then running the single-buyer algorithm is sufficient. On the other hand, in the cases where $C$ cannot be estimated, we observe an extra property of $R(p)$ that it is growing fast enough in $[0,C]$, leading to a “generalized half-concavity” property in $[0,p^{*}]$. This is sufficient to show the learnability of $R(p)$. Combining the two cases completes Step 1 of our algorithm.
Step 2: Binary Search with Prices $\hat{p}_{1},\cdots,\hat{p}_{n}$. The second step of the sub-routine is to update confidence intervals via binary searches. In this step, the main new challenge is the error from $\hat{p}_{1},\cdots\hat{p}_{n}$. For instance, while trying to update $[\ell_{i},r_{i}]$ with revenue function $R(p)=p\cdot(1-F(p))+C\cdot F(p)$, the constant $C$ comes from non-optimal prices $\hat{p}_{i+1},\cdots,\hat{p}_{n}$. Notice that $p^{*}_{i}$ is not the optimal price of $R(p)$. This creates an issue that comparing $R(\hat{p}_{i})$ with $R(p)$ for a price $p\in[\ell_{i},r_{i}]$ is not evident to show $p^{*}_{i}<p$ or $p^{*}_{i}>p$, preventing us from directly doing binary search.
To fix this issue, our main idea is to look at function $R^{*}(p):=p\cdot(1-F(p))+C^{*}\cdot F(p)$, where $C^{*}$ represents the revenue from the optimal prices $p^{*}_{i+1},\cdots p^{*}_{n}$, since $p^{*}_{i}$ is now the optimal price for $R^{*}(p)$. Although function $R^{*}(p)$ is not directly accessible, we can bound the difference between $C$ and $C^{*}$, and use the value of $R(p)$ to estimate $R^{*}(p)$. Running binary search with respect to this virtual function $R^{*}(p)$ completes Step 2 of our algorithm.
3.2 Notation
For the clarity of our algorithms, in this subsection we first introduce the following notations we will frequently use in our proofs.
  • $\delta$: We define $\delta=\frac{\epsilon}{100n^{2}}$ to be the scaled error parameter. $\delta$ also represents the unit length of those confidence intervals given by concentration inequalities.
  • $R(p_{1},\cdots,p_{n})$ and $p^{*}_{1},\cdots,p^{*}_{n}$: Recall that we defined $R(p_{1},\cdots,p_{n})$ to be the expected revenue when playing prices $(p_{1},\cdots,p_{n})$, and $(p^{*}_{1},\cdots,p^{*}_{n})$ represents the optimal prices that maximizes $R(p_{1},\cdots,p_{n})$.
  • $(\hat{p}_{1},\cdots,\hat{p}_{n})$: The approximately optimal price vector we find in first-step algorithm.
  • $S_{i}$: The suffix buyers $i,i+1,\cdots,n$. When finding $\hat{p}_{i-1}$, the performance of buyers $i,i+1,\cdots,n$ can be seen as a whole. $S_{i}$ represents the whole of this suffix starting from buyer $i$.
  • $r^{s}_{i}$: We define $r^{s}_{i}~{}:=~{}r_{i}-{T^{-100}}.$ 3.2 suggests that $r_{i}-{T^{-100}}$ is the highest price with regret guarantee. Therefore, $r^{s}_{i}$ represents this highest “safe” price for buyer $i$.
  • $P_{i}$: We define $P_{i}~{}:=~{}\prod_{j<i}F_{j}(r^{s}_{j}),$ where $F_{j}(\cdot)$ is the CDF of $\mathcal{D}_{j}$. $P_{i}$ is the probability that the item remains unsold for buyer $i$ when we set $p_{j}$ to be the highest safe price $r^{s}_{j}$ for all $j<i$, i.e., $P_{i}$ represents the maximum probability of testing buyer $i$.
  • $\mathsf{Loss}(S_{i})$: We define $\mathsf{Loss}(S_{i}):=R(1,\cdots,1,p^{*}_{i},\cdots,p^{*}_{n})-R(1,\cdots,1,\hat{p}_{i},\cdots,\hat{p}n)$, i.e., $\mathsf{Loss}(S_{i})$ represents the loss from $S_{i}$ when playing prices $\hat{p}_{i},\cdots,\hat{p}_{n}$.
  • $R_{i}(p)$: We define $R_{i}(p)~{}:=~{}R\big{(}r^{s}_{1},\cdot,r^{s}_{i-1},p_{i}=p,\hat{p}_{i+1},\cdots,\hat{p}_{n}\big{)}.$
    The intuition of function $R_{i}(p)$ is: When finding $\hat{p}_{i}$, naturally we hope to fix other prices to avoid interference from other buyers. For a buyer $j$ before $i$, we set $p_{j}$ to be $r^{s}_{j}$ to maximize the probability of testing $i$. For buyer $k$ from suffix $S_{i+1}$, we set $p_{k}$ to be the settled $\hat{p}_{k}$ to minimize the error from the suffix.
  • $\hat{p}^{*}_{i}$: We define $\hat{p}^{*}_{i}:=\arg\max_{p\in[\ell_{i},r_{i}]}R_{i}(p)$. It should be noted that $\hat{p}^{*}_{i}$ only represents the optimal price in the interval $[\ell_{i},r_{i}]$.
  • $R^{*}_{i}(p)$: We define $R^{*}_{i}(p)~{}:=~{}R(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i}=p,p^{*}_{i+1},\cdots,p^{*}_{n})$.
    The intuition of $R^{*}_{i}(p)$ is: notice that $\arg\max_{p}r^{*}_{i}(p)=p^{*}_{i}$. A key idea of our algorithm is to approach $p^{*}_{i}$ via estimating $R^{*}_{i}(p)$. It should be noted that $R^{*}_{i}(p)$ can be also written as $R^{*}_{i}(p)=R_{i}(p)+P_{i}\cdot F_{i}(p)\cdot\mathsf{Loss}(S_{i+1}).$ We will mainly use the second definition in our proofs.
3.3 $\widetilde{O}(\mathsf{poly}(n)\sqrt{T})$ Algorithm for $n$ Buyers: Proof of 3.1
We first prove our main 3.1 for BSPP assuming the following lemma for main sub-routine:
Lemma 3.2.
Given a BSPP instance with $R(p_{1},\cdots,p_{n})$ being the expected revenue function and $p^{*}_{1},\cdots,p^{*}_{n}$ being the optimal prices. Given error parameter $\epsilon\geq{1}/{\sqrt{T}}$ and intervals $[\ell_{1},r_{1}],\cdots,[\ell_{n},r_{n}]$ satisfying $p_{i}^{*}\in[\ell_{i},r_{i}]$ for all $i\in[n]$, there exists an algorithm that tests $O\big{(}\frac{n^{5}\log^{2}T}{\epsilon^{2}}\big{)}$ rounds, where each tested price vector $(p_{1},\cdots,p_{n})$ satisfies $p_{i}\in[\ell_{i},r_{i}-{T^{-100}}]$ for all $i\in[n]$, and outputs with probability $1-T^{-3}$ new confidence intervals $[\ell^{\prime}_{1},r^{\prime}_{1}]\subseteq[\ell_{1},r_{1}],\cdots,[\ell^{\prime}_{n},r^{\prime}_{n}]\subseteq[\ell_{n},r_{n}]$ satisfying
  • For $i\in[n]$, $p_{i}^{*}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$.
  • For every price vector $(p_{1},\cdots,p_{n})$ that satisfies $p_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}-{T^{-100}}]$ for all $i\in[n]$, we have
    $R(p^{*}_{1},\cdots,p^{*}_{n})~{}-~{}R(p_{1},\cdots,p_{n})~{}\leq~{}\epsilon.$
The algorithm in 3.2 uses some generalized ideas for single regular buyer. We defer the detailed algorithm and proofs to Section 3.4 and first give the $\widetilde{O}(\mathsf{poly}(n)\sqrt{T})$ regret algorithm assuming 3.2 holds.
Input: Hidden revenue function $R(p_{1},\cdots,p_{n})$, time horizon $T$
1 Let $\epsilon_{1}\leftarrow 1,[\ell^{(1)}_{j},r^{(1)}_{j}]\leftarrow[0,1]$ for all $j\in[n]$, and $i\leftarrow 1$.
2 while $\epsilon_{i}>\frac{n^{2.5}\log T}{\sqrt{T}}$ do
3       Run the algorithm described in 3.2 with $\epsilon_{i},[\ell^{(i)}_{1},r^{(i)}_{1}],\cdots,[\ell^{(i)}_{n},r^{(i)}_{n}]$ as input, and get $[\ell^{\prime}_{1},r^{\prime}_{1}],\cdots,[\ell^{\prime}_{n},r^{\prime}_{n}]$. Let $[\ell^{(i+1)}_{j},r^{(i+1)}_{j}]\leftarrow[\ell^{\prime}_{j},r^{\prime}_{j}]$ for all $j\in[n]$.
4       Let $\epsilon_{i+1}\leftarrow\frac{1}{2}\epsilon_{i}$ and $i\leftarrow i+1$.
5
Finish remaining rounds with any $(p_{1},\cdots,p_{n})$ satisfying $p_{j}\in[\ell^{(i)}_{j},r^{(i)}_{j}-{T^{-100}}]$ for all $j\in[n]$.
ALGORITHM 4 $O(n^{2.5}\sqrt{T}\log T)$ Regret Algorithm
Proof of 3.1.
We claim that the regret of Algorithm 4 is $O(n^{2.5}\sqrt{T}\log T)$ with probability $1-T^{-2}$. Algorithm 4 uses 3.2 for multiple times with a halving error parameter. Assume Algorithm 4 ends with $i=q+1$, i.e., the while loop runs $q$ times. Since we obtain $\epsilon_{i+1}=\epsilon_{i}/2$ and $\epsilon_{q}>\frac{n^{2.5}\log T}{\sqrt{T}}$ holds, there must be $q=O(\log T)$.
For $i\in[q]$, let $\mathsf{Alg}_{i}$ represent the corresponding algorithm we call when using 3.2 with $\epsilon_{i}$ and $\{[\ell^{(i)}_{j},r^{(i)}_{j}]\}$ as the input. To use 3.2 we need to verify that $p^{*}_{j}\in[\ell^{(i)}_{j},r^{(i)}_{j}]$ for all $i\in[q],j\in[n]$. The condition $p^{*}_{j}\in[\ell^{(1)}_{j},r^{(1)}_{j}]=[0,1]$ clearly holds. For $i=2,3,\cdots,q$, the condition $p^{*}_{j}\in[\ell^{(i)}_{j},r^{(i)}_{j}]$ is guaranteed by 3.2 when calling $\mathsf{Alg}_{i-1}$. The failing probability of 3.2 is $T^{-3}$. By the union bound, with probability $1-T^{-2}$, 3.2 always holds.
Now we give the regret of Algorithm 4. 3.2 guarantees that for $i\geq 2$, every price vector we test in $\mathsf{Alg}_{i}$ has regret at most $\epsilon_{i-1}$. Hence, the total regret of Algorithm 4 can be bounded by
$1\cdot O\big{(}\epsilon_{1}^{-2}n^{5}\log^{2}T\big{)}~{}+~{}\sum_{i=2}^{q}\epsilon_{i-1}\cdot O\big{(}\epsilon^{-2}_{i}n^{5}\log^{2}T\big{)}~{}+~{}T\cdot\epsilon_{q}~{}=~{}O(n^{5}\log^{2}T+n^{2.5}\sqrt{T}\log T).$
In this paper, we consider the setting that $n$ is a fixed parameter while $T$ goes to infinity, and further assume that $n^{5}=o(\sqrt{T})$, otherwise an $O(\mathsf{poly}(n))$ regret algorithm is trivial. Then, $O(n^{5}\log^{2}T)$ can be upper bounded by $o(n^{2.5}\sqrt{T}\log T)$ and the regret bound of Algorithm 4 is $O(n^{2.5}\sqrt{T}\log T)$.
Finally, we also need to check that Algorithm 4 uses no more than $T$ rounds. 3.2 suggests that $\mathsf{Alg}_{i}$ runs $O(\epsilon^{-2}_{i}n^{5}\log^{2}T)$ rounds. So the total number of rounds in the while loop is $\sum_{i\in[q]}O(\epsilon^{-2}_{i}n^{5}\log^{2}T)~{}=~{}O(T).$ Therefore, Algorithm 4 is feasible. ∎
3.4 Two-Steps Sub-Routine: Proof of 3.2
In this section, we present the main sub-routine for sequential buyers and prove 3.2. Our algorithm uses a two-steps algorithm similar to the single-buyer setting to update all intervals: the first step is to find a group of approximately optimal prices $(\hat{p}_{1},\cdots,\hat{p}_{n})$, and the second step is to use $(\hat{p}_{1},\cdots,\hat{p}_{n})$ as a benchmark to update the upper and lower bounds via binary search. We show that in general this approach works after overcoming additional challenges. Comparing to the single-buyer algorithm, the main extra challenge is that the half-concavity does not always hold for sequential buyers. To fix this issue, we define a generalized version of half-concavity and give a new algorithm that works for this generalized half-concavity.
As discussed in Section 3.1, the main sub-routine defined in 3.2 contains two steps: the first step is to find a group of approximately optimal prices $(\hat{p}_{1},\cdots,\hat{p}_{n})$, and the second step is to use $(\hat{p}_{1},\cdots,\hat{p}_{n})$ as a benchmark to update the upper and lower bounds via binary search. We first give corresponding key lemmas for these two steps.
Lemma 3.3 (Step 1).
Given $\delta:=\frac{\epsilon}{100n^{2}}\geq\frac{1}{\sqrt{T}}$, intervals $[\ell_{1},r_{1}],\cdots,[\ell_{n},r_{n}]$ that satisfies $p^{*}_{i}\in[\ell_{i},r_{i}]$ for all $i\in[n]$, there exists an algorithm that tests $O(\frac{n\log^{2}T}{\delta^{2}})$ rounds with price vectors $(p_{1},\cdots,p_{n})$ that satisfy $p_{i}\in[\ell_{i},r^{s}_{i}]$ for all $i\in[n]$ and outputs $(\hat{p}_{1},\cdots,\hat{p}_{n})$. With probability $1-T^{-4}$, we have the following for all $i\in[n]$:
  • $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$.
  • $P_{i}\cdot\mathsf{Loss}_{i}\leq 5(n-i+1)\delta$.
Lemma 3.4 (Step 2).
Given $\delta:=\frac{\epsilon}{100n^{2}}\geq\frac{1}{\sqrt{T}}$, intervals $[\ell_{1},r_{1}],\cdots,[\ell_{n},r_{n}]$ that satisfies $p^{*}_{i}\in[\ell_{i},r_{i}]$ for all $i\in[n]$, and near-optimal prices $\hat{p}_{1},\cdots,\hat{p}_{n}$ that satisfy the conditions in 3.3, there exists an algorithm that tests $O(\frac{n\log^{2}T}{\delta^{2}})$ rounds with price vectors $(p_{1},\cdots,p_{n})$ that satisfy $p_{i}\in[\ell_{i},r^{s}_{i}]$ for all $i\in[n]$ and outputs $[\ell^{\prime}_{1},r^{\prime}_{1}],\cdots,[\ell^{\prime}_{n},r^{\prime}_{n}]$. With probability $1-T^{-4}$, we have the following:
  • $[\ell^{\prime}_{i},r^{\prime}_{i}]\subseteq[\ell_{i},r_{i}]$.
  • $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$.
  • For every price vector $(p_{1},\cdots,p_{n})$ that satisfies $p_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}-{T^{-100}}]$ for all $i\in[n]$, we have
    $R(p^{*}_{1},\cdots,p^{*}_{n})~{}-~{}R(p_{1},\cdots,p_{n})~{}\leq~{}\epsilon.$
Combining 3.3 and 3.4 immediately proves 3.2: we run the algorithms defined in 3.3 and 3.4 sequentially. Both algorithms use $O(\delta^{-2}\cdot n\log T)=O(\epsilon^{-2}\cdot n^{5}\log T)$ rounds, so the constraint on number of rounds in 3.2 is satisfied. The feasibility of each tested price vector is guaranteed by both 3.3 and 3.4. Finally, for the required conditions of the output, each algorithm succeeds with probability $1-T^{-4}$. By the union bound, with probability $1-2T^{-4}>1-T^{-3}$, both algorithms succeed simultaneously. Then, the statements in 3.2 is guaranteed by 3.4.
It only remains to prove 3.3 and 3.4. Next, we prove 3.3 in Section 3.5, and 3.4 in Section 3.6.
3.5 Step 1: Finding Near-Optimal $(\hat{p}_{1},\cdots,\hat{p}_{n})$ to Prove 3.3
In this section, we give the algorithm defined in 3.3. Our algorithm is based on a key sub-algorithm that finds $\hat{p}_{i}$ such that $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$. The following lemma states the existence of such a sub-algorithm:
Lemma 3.5.
There exists a sub-algorithm $\mathsf{SubAlg}$ that takes hidden function $R_{i}(p)$, interval $[\ell_{i},r_{i}]$, and parameter $\delta$ as input, tests $O(\delta^{-2}\log^{2}T)$ rounds with price vectors $(r^{s}_{1},\cdots,r^{s}_{i-1},p\in[\ell_{i},r^{s}_{i}],$ $\hat{p}_{i+1},\cdots,\hat{p}_{n})$, and outputs $\hat{p}_{i}\in[\ell_{i},r^{s}_{i}]$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ with probability $1-T^{-5}$.
The design of $\mathsf{SubAlg}$ requires the idea of generalized half-concavity discussed in Section 3.1. Since the proof requires some involved technical details, we defer it separately to Section 3.7. With the help of 3.5, we give our algorithm and the proof of 3.3.
Input: Hidden Revenue Function $R(p_{1},\cdots,p_{n})$, Intervals $[\ell_{1},r_{1}],\cdots,[\ell_{n},r_{n}]$, Error Parameter $\epsilon$
1 Let $\delta=\frac{\epsilon}{100n^{2}}$ be the scaled error parameter.
2 for $i=n\to 1$ do
3       Call $\mathsf{SubAlg}$ with input $R_{i}(p),[\ell_{i},r_{i}],\delta$ and get $\hat{p}_{i}$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$.
Output: $\hat{p}_{1},\cdots,\hat{p}_{n}$
ALGORITHM 5 Finding $\hat{p}_{1},\cdots,\hat{p}_{n}$
Proof of 3.3.
We show that Algorithm 5 is the desired algorithm for 3.3. 3.5 guarantees that calling $\mathsf{SubAlg}$ one time runs $O(\frac{\log^{2}T}{\delta^{2}})$ rounds. Therefore, Algorithm 5 tests $O(\frac{n\log^{2}T}{\delta^{2}})$ rounds in total. For the succeed probability, Algorithm 5 calls $\mathsf{SubAlg}$ $n$ times. Each call fails with probability $T^{-5}$. By the the union bound [^2] Recall that we assumed $n=o(T)$., with probability $1-T^{-4}$, all calls of $\mathsf{SubAlg}$ in Algorithm 5 succeed. Then, the constraints $p_{i}\in[\ell_{i},r_{i}-{T^{-100}}]$ and $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ are directly guaranteed by 3.5.
It only remains to prove $P_{i}\cdot\mathsf{Loss}(S_{i})\leq 5(n-i+1)\delta$. Now we prove this statement assuming $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ holds.
We do the prove via induction from $i=n$ to $i=1$. The base case is $i=n$. We have
$P_{n}\cdot\mathsf{Loss}(S_{n})~{}=~{}R_{n}(p^{*}_{n})-R_{n}(\hat{p}_{n})~{}=~{}R_{n}(\hat{p}^{*}_{n})-R_{n}(\hat{p}_{n})\leq 4\delta<5\delta.$
For the induction step, we prove $P_{i}\cdot\mathsf{Loss}(S_{i})\leq 5(n-i+1)\delta$ under the induction hypothesis $P_{i+1}\cdot\mathsf{Loss}(S_{i+1})\leq 5(n-i)\delta$. We prove the induction step by discussing the following two cases.
Case 1: $p^{*}_{i}\in[\ell_{i},r^{s}_{i}]$. In this case, we have
$P_{i}\cdot\mathsf{Loss}(S_{i})~{}=~{}$ $R(r^{s}_{1},\cdots,r^{s}_{i-1},p^{*}_{i},p^{*}_{i+1}\cdots,p^{*}_{n})-R(r^{s}_{1},\cdots,r^{s}_{i-1},\hat{p}_{i},\hat{p}_{i+1}\cdots,\hat{p}_{n})$
[…]
$+R(r^{s}_{1},\cdots,r^{s}_{i-1},p^{*}_{i},\hat{p}_{i+1}\cdots,\hat{p}_{n})-R(r^{s}_{1},\cdots,r^{s}_{i-1},\hat{p}_{i},\hat{p}_{i+1}\cdots,\hat{p}_{n})$
[…]
Case 2: $p^{*}_{i}\in[r^{s}_{i},r_{i}]$. For this case, we have
$P_{i}\cdot\mathsf{Loss}(S_{i})~{}=~{}$ $R(r^{s}_{1},\cdots,r^{s}_{i-1},p^{*}_{i},p^{*}_{i+1}\cdots,p^{*}_{n})-R(r^{s}_{1},\cdots,r^{s}_{i-1},\hat{p}_{i},\hat{p}_{i+1}\cdots,\hat{p}_{n})$
[…]
$+R(r^{s}_{1},\cdots,r^{s}_{i-1},r^{s}_{i},\hat{p}_{i+1}\cdots,\hat{p}_{n})-R(r^{s}_{1},\cdots,r^{s}_{i-1},\hat{p}_{i},\hat{p}_{i+1}\cdots,\hat{p}_{n})+\delta$
[…]
where the first inequality is from the following fact: let $C:=R(1,\cdots,1,p_{i}=1,p^{*}_{i+1},\cdots,p^{*}_{n})$ be the optimal revenue from suffix $S_{i}$, we have
$R(r^{s}_{1},\cdots,r^{s}_{i-1},p^{*}_{i},p^{*}_{i+1}\cdots,p^{*}_{n})-R(r^{s}_{1},\cdots,r^{s}_{i-1},r^{s}_{i},p^{*}_{i+1}\cdots,p^{*}_{n})$
[…]
where the second inequality is from the fact that $p^{*}_{i}\geq C$, because $p^{*}_{i}$ is the maximizer of function $R^{*}_{i}(p)$, and $R^{*}_{i}(p^{*}_{i})-R^{*}_{i}(C)=P_{i}\cdot(1-F_{i}(p^{*}_{i}))\cdot(p^{*}_{i}-C)\geq 0$ implies $p^{*}_{i}\geq C$.
Therefore, we have $P_{i}\mathsf{Loss}(S_{i})\leq 5(n-i+1)\delta$ in both cases, which finishes the induction step. ∎
3.6 Step 2: Determining New Intervals via Binary Search to Prove 3.4
The second step of our subroutine is to use prices $\hat{p}_{1},\cdots,\hat{p}_{n}$ as a benchmark to determine new confidence intervals. Comparing to the single-buyer setting, there exists an extra challenge for sequential buyers: We only have access to the function $R_{i}(p)$. However, the optimal price for $R_{i}(p)$ is $\hat{p}^{*}_{i}$, instead of the desired $p^{*}_{i}$. At a first glance, it is unreasonable to determine the confidence interval for $p^{*}_{i}$ only with access to function $R_{i}(p)$. To fix this issue, we introduce
$R^{*}_{i}(p)~{}:=~{}R(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i}=p,p^{*}_{i+1},\cdots,p^{*}_{n})~{}=~{}R_{i}(p)+P_{i}\cdot F_{i}(p)\cdot\mathsf{Loss}(S_{i+1}).$
Then, $R^{*}_{i}(p)$ acts as a bridge between $R_{i}(p)$ and $p^{*}_{i}$, because we have $p^{*}_{i}=\arg\max_{p}R^{*}_{i}(p)$, and the difference between $R_{i}(p)$ and $R^{*}_{i}(p)$ can be bounded via the following claim:
Claim 3.6.
If 3.3 holds, we have $R_{i}(p)\leq R^{*}_{i}(p)\leq R_{i}(p)+5(n-i)\delta$ for all $i\in[n],p\in[\ell_{i},r^{s}_{i}]$.
Proof.
$R_{i}(p)\leq R^{*}_{i}(p)$ holds simply because $\mathsf{Loss}(S_{i+1})\geq 0$. For $R^{*}_{i}(p)\leq R_{i}(p)+5(n-i)\delta$, we have $P_{i}\cdot\mathsf{Loss}(S_{i})\leq 5(n-i+1)\delta$ when 3.3 holds. Notice that $P_{i}\cdot F_{i}(p)\mathsf{Loss}(S_{i+1})\leq P_{i+1}\mathsf{Loss}(S_{i+1})\leq 5(n-i)\delta$ when $p\in[\ell_{i},r^{s}_{i}]$. This gives the inequality $R^{*}_{i}(p)\leq R_{i}(p)+5(n-i)\delta$. ∎
With 3.6, the idea of our binary search algorithm is clear: We first set $R_{i}(\hat{p}_{i})$ be the benchmark. For a price $p$, we can estimate $R_{i}(p)$ and compare $R_{i}(p)+5(n-i)\delta$ with $R_{i}(\hat{p}_{i})$. If $R_{i}(p)+5(n-i)\delta<R_{i}(\hat{p}_{i})$, it is evident to say $R^{*}_{i}(p)<R^{*}_{i}(\hat{p}_{i})$, and therefore $p$ can’t be $p^{*}_{i}$. Otherwise, we keep $p$ as a candidate of $p^{*}_{i}$. Since function $R_{i}(p)$ is single-peaked, the experiment above can be done via two binary search algorithms, and extra steps are needed to take care of the errors from the uncertainty - we only have estimates $\hat{R}_{i}(p)$ and $\hat{R}_{i}(\hat{p}_{i})$, instead of $R_{i}(p)$ and $R_{i}(\hat{p}_{i})$. Now, we state the following lemma that achieves our idea:
Lemma 3.7.
Assume 3.3 holds. Given hidden function $R_{i}(p)$, interval $[\ell_{i},r_{i}]\ni p^{*}_{i}$, scaled $\delta:=\frac{\epsilon}{100n^{2}}>1/\sqrt{T}$, price $\hat{p}_{i}$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$, there exists a binary search algorithm that runs $O(\frac{\log^{2}T}{\delta^{2}})$ rounds with price vectors $(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i}\in[\ell_{i},r^{s}_{i}],\hat{p}_{i+1},\cdots,\hat{p}_{n})$ and outputs with probability $1-T^{-5}$ $[\ell^{\prime}_{i},r^{\prime}_{i}]\subseteq[\ell_{i},r_{i}]$ that satisfies:
  • $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$.
  • For $p\in[\ell^{\prime}_{i},r^{\prime}_{i}-{T^{-100}}]$, $R_{i}(\hat{p}^{*}_{i})-R_{i}(p)\leq 5(n-i)\delta+15\delta$.
To prove 3.7, we first give the following Algorithm 6, and then show Algorithm 6 is the desired algorithm for 3.7.
Input: Hidden Revenue Function $R_{i}(p)$, Interval $[\ell_{i},r_{i}]$, Error Parameter $\delta$
Fix $p_{j}=r^{s}_{j}$ for $j<i$ and $p_{j}=\hat{p}_{j}$ for $j>i$. // First binary search to determine $\ell^{\prime}_{i}$
1 Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p_{i}=\hat{p}_{i}$. Let $\hat{R}_{i}(\hat{p}_{i})$ be the average reward.
2 Let $\ell_{b1}\leftarrow\ell_{i},r_{b1}\leftarrow\hat{p}_{i}$.
3 while $r_{b1}-\ell_{b1}\geq{T^{-100}}$ do
4       Let $m=\frac{\ell_{b1}+r_{b1}}{2}$.
5       Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p_{i}=m$. Let $\hat{R}_{i}(m)$ be the average reward.
6       if $\hat{R}_{i}(m)<\hat{R}_{i}(\hat{p}_{i})-2\delta-5(n-i)\delta$ then Update $r_{b1}\leftarrow m$ else Update $\ell_{b1}\leftarrow m$;
7
8Test $p_{i}=\ell_{b1}$ and $p_{i}=r_{b1}$ each for $C\cdot\frac{\log T}{\delta^{2}}$ rounds. Let $\hat{R}_{i}(\ell_{b1}),\hat{R}_{i}(r_{b1})$ be the average reward.
9 if $\hat{R}_{i}(\ell_{b1})<\hat{R}_{i}(r_{b1})-3\delta$ then Let $\ell^{\prime}_{i}\leftarrow r_{b1}$ else Let $\ell^{\prime}_{i}\leftarrow\ell_{b1}$;
  // Second binary search to determine $r^{\prime}_{i}$
10 Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p_{i}=r_{i}-{T^{-100}}$. Let $\hat{R}_{i}(r_{i}-{T^{-100}})$ be the average reward.
11 if $\hat{R}_{i}(r-{T^{-100}})\geq\hat{R}_{i}(\hat{p}_{i})-2\delta$ then
       // Special Case: original upper bound is good
12       Let $r^{\prime}_{i}\leftarrow r_{i}$
13else
14       Let $\ell_{b2}\leftarrow\hat{p}_{i},r_{b2}\leftarrow r_{i}-{T^{-100}}$.
15       while $r_{b2}-\ell_{b2}\geq{T^{-100}}$ do
16             Let $m=\frac{\ell_{b2}+r_{b2}}{2}$.
17             Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p_{i}=m$. Let $\hat{R}_{i}(m)$ be the average reward.
18             if $\hat{R}_{i}(m)<\hat{R}_{i}(\hat{p})-2\delta-5(n-i)\delta$ then Update $r_{b2}\leftarrow m$ else Update $\ell_{b2}\leftarrow m$;
19
20      Let $r^{\prime}_{i}\leftarrow r_{b2}$.
Output: $[\ell^{\prime}_{i},r^{\prime}_{i}]$.
ALGORITHM 6 Getting $[\ell^{\prime}_{i},r^{\prime}_{i}]$ via Binary Search
Proof of 3.7.
We show Algorithm 6 satisfies the statements in 3.7. Our proof starts from verifying Algorithm 6 runs $O(\frac{\log^{2}T}{\delta^{2}})$ rounds with $p_{i}\in[\ell_{i},r_{i}-{T^{-100}}]$. For the number of tested rounds, Algorithm 6 runs two standard binary searches. In one binary search, the subroutine tests one price with $C\cdot\frac{\log T}{\delta^{2}}$ rounds, and the binary search stops when the length of the interval is less than ${T^{-100}}$. Therefore, The total number of rounds is $C\cdot\frac{\log T}{\delta^{2}}\cdot O(\log T^{100})=O(\frac{\log^{2}T}{\epsilon^{2}})$. For the range constraint of all tested prices, Algorithm 6 directly guarantees that every tested price falls in $[\ell_{i},r_{i}-{T^{-100}}]$.
Now we move to the two main statements. We first give the following claim:
Claim 3.8.
In Algorithm 6, $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta|$ simultaneously holds for every tested price $p$ with probability at least $1-T^{-5}$.
Proof.
For a single tested price $p$, $\hat{R}(p)$ is estimated by calculating the average of $N=C\cdot\frac{\log T}{\delta^{2}}$ samples. By Hoeffding’s Inequality,
$\mbox{\rm\bf Pr}\left[|\hat{R}_{i}(p)-R_{i}(p)|>\delta\right]~{}\leq~{}2\exp\left(-2N\delta^{2}\right)~{}=~{}2T^{-2C}~{}<~{}T^{-6}.$
The last inequality holds when $C$ is a constant greater than $4$. Then, we have $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ holds with probability $1-T^{-6}$ for a single tested price $p$.
Notice that Algorithm 6 can’t test more than $T$ prices. By the union bound, $|\hat{R}(p)-R(p)|\leq\delta$ simultaneously holds for all tested prices with probability $1-T^{-5}$. ∎
We will give the remaining parts of the proof assuming $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ always holds.
Now we show $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$. The key idea is to use 3.6. For clarity we first restate 3.6: See 3.6
Our goal is to use binary search to determine $I^{\prime}:=\{p\in[\ell^{\prime}_{i},r^{\prime}_{i}]:R_{i}(p)+5(n-i)\delta\geq R_{i}(\hat{p}_{i})\}$. For $p\in[\ell^{\prime}_{i},r^{\prime}_{i}]\setminus I^{\prime}$, there must be
$R^{*}_{i}(p)~{}\leq~{}R_{i}(p)+5(n-i)\delta~{}<~{}R_{i}(\hat{p}_{i})~{}\leq~{}R^{*}_{i}(\hat{p}_{i})~{}\leq R^{*}_{i}(p^{*}_{i}).$
Recall that $p^{*}_{i}$ is the optimal price for $R^{*}_{i}(p)$. Therefore, the price $p\in[\ell^{\prime}_{i},r^{\prime}_{i}]\setminus I^{\prime}$ can’t be $p^{*}_{i}$, i.e., $p^{*}_{i}$ falls in $I^{\prime}$.
Now we prove $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$. We first show that $p^{*}_{i}\geq\ell_{b1}$. Here, $\ell_{b1}$ refers to the value after the while loop is finished. $p^{*}_{i}\geq\ell_{b1}$ holds because $\ell_{b1}$ is updated to be at least $m$ when we observe $\hat{R}_{i}(m)<\hat{R}_{i}(\hat{p}_{i})-2\delta-5(n-i)\delta$. Combining this with 3.8 implies $R_{i}(m)<R_{i}(\hat{p}_{i})-5(n-i)\delta$, i.e., $m\notin I^{\prime}$. Since $R_{i}(p)$ is single-peaked, it’s sufficient to update $\ell_{b1}$ to at least $m$ and $I^{\prime}\subseteq[\ell_{b1},1]$ holds. Symmetrically, a similar argument gives $p^{*}_{i}\leq r_{b2}=r^{\prime}_{i}$.
It remains to show $p^{*}_{i}\geq\ell^{\prime}_{i}$. $\ell^{\prime}_{i}$ and $\ell_{b1}$ are different when we have $\hat{R}(\ell_{b1})<\hat{R}(r_{b1})-3\delta$. Combining this inequality with 3.8 gives $R(r_{b1})-R(\ell_{b1})>\delta>r_{b1}-\ell_{b1}$. This is in contrast to the constrained Lipschitzness from 3.9. Then, there must be $r_{b1}<\mathsf{Rev}(S_{i+1})<\mathsf{Opt}(S_{i+1})\leq p^{*}_{i}$, where the last inequality $p^{*}_{i}\geq\mathsf{Opt}(S_{i+1})$ is implied by the fact that $R^{*}_{i}(p^{*}_{i})-R^{*}_{i}(\mathsf{Opt}(S_{i+1}))=P_{i}\cdot(1-F_{i}(p^{*}_{i}))\cdot(p^{*}_{i}-\mathsf{Opt}(S_{i+1}))\geq 0$. Therefore, it’s feasible to update $\ell^{\prime}_{i}$ to $r_{b1}$ in this case.
Finally, we show $R_{i}(\hat{p}^{*}_{i})-R_{i}(p)\leq 5(n-i)\delta+15\delta$ for all $p\in[\ell^{\prime}_{i},r^{\prime}_{i}-{T^{-100}}]$. Function $R(p)$ is single-peaked, so it is sufficient to prove $R_{i}(\hat{p}^{*}_{i})-R_{i}(\ell^{\prime}_{i})\leq 5(n-i)\delta+11\delta$ and $R_{i}(\hat{p}^{*}_{i})-R_{i}(r^{\prime}_{i}-{T^{-100}})\leq 5(n-i)\delta+11\delta$. Then, combining these two inequalities with the assumption $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ gives the desired statement.
We first prove that $R_{i}(\hat{p}_{i})-R(\ell^{\prime}_{i})\leq 5(n-i)\delta+11\delta$. The binary search subroutine guarantees that $\hat{R}_{i}(r_{b1})\geq\hat{R}_{i}(\hat{p}_{i})-2\delta-5(n-i)\delta$ after the binary search loop is finished. Then, 3.8 ensures that $R_{i}(\hat{p}_{i})-R_{i}(r_{b1})\leq 5(n-i)\delta+4\delta$. Then, if $\hat{R}_{i}(\ell_{b1})<\hat{R}_{i}(r_{b1})-3\delta$, we set $\ell^{\prime}_{i}$ to be $r_{b1}$. Otherwise, we have $\hat{R}_{i}(\ell_{b1})\geq\hat{R}_{i}(r_{b1})-3\delta$ and $\ell^{\prime}_{i}$ is set to be $\ell_{b1}$. Then,
$R_{i}(\hat{p}_{i})-R_{i}(\ell^{\prime}_{i})~{}\leq$ $~{}R_{i}(r_{b1})-R_{i}(\ell_{b1})+5(n-i)\delta+4\delta$
[…]
Therefore, in both cases $R_{i}(\hat{p}_{i})-R(\ell^{\prime}_{i})\leq 11(n-i)\delta+6\delta$ holds.
For the inequality $R_{i}(\hat{p}_{i})-R_{i}(r^{\prime}_{i}-{T^{-100}})\leq 5(n-i)\delta+11\delta$, a symmetric proof for the second binary search subroutine gives $R_{i}(\hat{p})-R_{i}(\ell_{b2})\leq 5(n-i)\delta+4\delta$, while $r^{\prime}_{i}$ is set to be $r_{b2}<\ell_{b2}+{T^{-100}}$. If $\hat{p}^{*}_{i}<r^{\prime}_{i}-{T^{-100}}$, the single-peakedness of $R(p)$ gives
$R_{i}(r^{\prime}_{i}-{T^{-100}})~{}\geq~{}R_{i}(\ell_{b2})~{}\geq~{}R_{i}(\hat{p})-5(n-i)\delta-4\delta.$
Otherwise, the constrained Lipschitzness of $R_{i}(p)$ gives
$R_{i}(r^{\prime}_{i}-{T^{-100}})~{}\geq~{}R_{i}(\ell_{b2})-(\ell_{b2}-r_{b2}+{T^{-100}})~{}\geq~{}R_{i}(\hat{p})-5(n-i)\delta-4\delta-{T^{-100}}>R_{i}(\hat{p})-5(n-i)\delta-5\delta.\qed$
Finally, we end this section proving 3.4 via 3.7.
Proof of 3.4.
The algorithm for 3.4 is to use 3.7 $n$ times and get $n$ new confidence intervals. 3.7 succeeds with probability $1-T^{-5}$. By the union bound, with probability $1-n\cdot T^{-5}>1-T^{-4}$, our algorithm for 3.4 succeeds, and the conditions $[\ell^{\prime}_{i}r^{\prime}_{i}]\subseteq[\ell_{i},r_{i}]$ and $p^{*}_{i}\in[\ell^{\prime}_{i},r^{\prime}_{i}]$ are guaranteed by 3.7. It only remains to show that each price vector in the new intervals has regret no more than $\epsilon$.
We instead prove the following statement via induction, assuming the statements in 3.3 and 3.7 hold: For $p_{i},\cdots,p_{n}$ that satisfies $p_{j}\in[\ell^{\prime}_{j},r^{\prime}_{j}-{T^{-100}}]$ for all $j\in\{i,\cdots,n\}$, we have
$R(r^{s}_{1},\cdots,r^{s}_{i-1},p^{*}_{i},\cdots,p^{*}_{n})~{}-~{}R(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i},\cdots,p_{n})~{}\leq~{}50(n-i+1)^{2}\delta.$
The base case is $i=n$. 3.7 gives
$R(r^{s}_{1},\cdots,r^{s}_{n-1},p^{*}_{n})~{}-~{}R(r^{s}_{1},\cdots,r^{s}_{n-1},p_{n})~{}\leq~{}15\delta<50\delta.$
For the induction step, we hope to show
$R(r^{s}_{1},\cdots,r^{s}_{i-1},p^{*}_{i},\cdots,p^{*}_{n})~{}-~{}R(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i},\cdots,p_{n})~{}\leq~{}50(n-i+1)^{2}\delta$
under the induction hypothesis
$R(r^{s}_{1},\cdots,r^{s}_{i-1},r^{s}_{i},p^{*}_{i+1},\cdots,p^{*}_{n})~{}-~{}R(r^{s}_{1},\cdots,r^{s}_{i-1},r^{s}_{i},p_{i+1},\cdots,p_{n})~{}\leq~{}50(n-i)^{2}\delta.$
We first observe that in the induction hypothesis, if we change the price of buyer $i$ from $r^{s}_{i}$ to $p_{i}$, it only reduces the probability of incurring a loss. Therefore, the induction hypothesis also gives
$R(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i},p^{*}_{i+1},\cdots,p^{*}_{n})~{}-~{}R(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i},p_{i+1},\cdots,p_{n})~{}\leq~{}50(n-i)^{2}\delta.$
Now we finish the induction step:
$R(r^{s}_{1},\cdots,r^{s}_{i-1},p^{*}_{i},p^{*}_{i+1},\cdots,p^{*}_{n})~{}\leq$ $~{}R(r^{s}_{1},\cdots,r^{s}_{i-1},\hat{p}_{i},\hat{p}_{i+1},\cdots,\hat{p}_{n})+5(n-i+1)\delta$
[…]
$+10(n-i+1)\delta+14\delta+50(n-i)^{2}\delta$
[…]
Here, the first inequality is from the inequality $P_{i}\mathsf{Loss}(S_{i})\leq 5(n-i+1)\delta$ in 3.3. The second inequality is from 3.3 that guarantees $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$. The third inequality is from 3.7 that guarantees $R_{i}(\hat{p}^{*}_{i})-R_{i}(p_{i})\leq 5(n-i)\delta+15\delta$. The fourth inequality is from the fact that prices $p^{*}_{i+1},\cdots,p^{*}_{n}$ are optimal for suffix $S_{i+1}$. The fifth inequality is from the corollary of the induction hypothesis. Comparing the start and the end of the above inequality finishes the induction step.
Finally, taking $i=1$ with the fact that $\delta=\frac{\epsilon}{100n^{2}}$ finishes the proof of 3.4. ∎
3.7 Sub-Algorithm $\mathsf{SubAlg}$: Proof of Lemma 3.5
We now introduce the sub-algorithm $\mathsf{SubAlg}$ that finds $\hat{p}_{i}$ satisfying $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$, and prove the missing 3.5 that will complete the proof of Step 1. We restate the lemma for convenience.
See 3.5
We start from introducing the following new notations that we need in this subsection:
  • $\mathsf{Pre}(S_{i})$: We define $\mathsf{Pre}(S_{i}):=R(r^{s}_{1},\cdots,r^{s}_{i-1},1,\cdots,1)$ to be the expected revenue from the buyers before $S_{i}$ with prices $r^{s}_{1},\cdots,r^{s}_{i-1}$.
  • $\mathsf{Opt}(S_{i})$: We define $\mathsf{Opt}(S_{i}):=R(1,\cdots,1,p^{*}_{i},\cdots,p^{*}_{n})$ to be the optimal revenue from $S_{i}$.
  • $\mathsf{Rev}(S_{i})$: We define $\mathsf{Rev}(S_{i}):=R(1,\cdots,1,\hat{p}_{i},\cdots,\hat{p}n)$ to be the near-optimal revenue from $S_{i}$ with near-optimal prices $\hat{p}_{i},\cdots,\hat{p}_{n}$.
With the help of these new notations, we can rewrite $R_{i}(p)$ as
$R_{i}(p)~{}=~{}\mathsf{Pre}(S_{i})+P_{i}(p\cdot(1-F_{i}(p))+\mathsf{Rev}(S_{i+1})\cdot F_{i}(p)).$
Comparing to the revenue function of a single buyer, function $R_{i}(p)$ has an extra $\mathsf{Rev}(S_{i+1})\cdot F_{i}(p)$ term, and therefore half-concavity is not directly applicable. The good news is that $R_{i}(p)$ is still half-concave in part of the interval:
Claim 3.9.
Function $R_{i}(p)$ satisfies the following properties:
  • $R_{i}(p)$ is single-peaked in $[0,1]$ with $\arg\max_{p\in[0,1]}\geq\mathsf{Rev}(S_{i+1})$.
  • $R_{i}(p)$ has constrained Lipschitzness: For $b\in[\mathsf{Rev}(S_{i+1}),1]$ and $a\in[0,b]$, we have $R(b)-R(a)\leq P_{i}(b-a)$.
  • $R_{i}(p)$ is half-concave in $[\mathsf{Rev}(S_{i+1}),1]$.
Proof.
By the definition of $R_{i}(p)$, we have $R_{i}(p)=P_{i}(p\cdot(1-F_{i}(p)+C_{i}\cdot F_{i}(p))$, where $P_{i}:=\prod_{j<i}F_{j}(r_{j})$ is the probability of reaching buyer $i$. Now we prove the three properties one-by-one: (i) Single-peakedness: Let $R_{q}(q)$ be the revenue function in the quantile space that correspond to $R_{i}(p)$. By definition we have $R_{q}(q)=P_{i}(q^{-1}(q)\cdot(1-q)+C_{i}\cdot q)$. The key observation is that comparing to the original concave revenue function $q^{-1}(q)\cdot q$, $R_{q}(q)$ has an extra linear term, so it still remains concave. Therefore, function $R_{i}(p)$ is single-peaked in the value space.
We also need to show that $\arg\max_{p}R_{i}(p)\geq\mathsf{Rev}(S_{i+1})$. This is true because for all $p\in[0,\mathsf{Rev}(S_{i+1})]$,
$R_{i}(p)~{}=$ $~{}\mathsf{Pre}(S_{i})+P_{i}(p\cdot(1-F_{i}(p))+\mathsf{Rev}(S_{i+1})\cdot F_{i}(p))$
[…]
The above inequality shows $R_{i}(\mathsf{Rev}(S_{i+1}))$ is not worse than any price in $[0,\mathsf{Rev}(S_{i+1})]$. Therefore, there must be $\arg\max_{p}R_{i}(p)\geq\mathsf{Rev}(S_{i+1})$.
(ii) Constrained Lipschitzness: For $b\geq\mathsf{Rev}(S_{i+1})$ and $a\leq b$, we have
$R_{i}(b)-R_{i}(a)~{}=~{}P_{i}((b-a)-(b-\mathsf{Rev}(S_{i+1}))F_{i}(b)+(a-\mathsf{Rev}(S_{i+1}))F_{i}(a))~{}\leq~{}P_{i}(b-a).$
Here, the inequality is true because we have $b-\mathsf{Rev}(S_{i+1})\geq a-\mathsf{Rev}(S_{i+1})$, $b-\mathsf{Rev}(S_{i+1})\geq 0$, and $F_{i}(b)\geq F_{i}(a)\geq 0$. Then, for quantities $A,B,C,D$ satisfying $A\geq 0,A\geq B,C\geq D\geq 0$, it’s easy to verify that $A\cdot C\geq B\cdot D$.
(iii) Half-Concavity in $[\mathsf{Rev}(S_{i+1}),1]$: The single-peakedness and the Lipschitzness required by the definition of half-concavity are both proved. Let $p^{*}=\arg\!\max_{p\in[0,1]}R_{i}(p)$. We proved that $p^{*}\geq\mathsf{Rev}(S_{i+1})$, then it only remains to show $R_{i}(p)$ is concave in $[\mathsf{Rev}(S_{i+1}),p^{*}]$.
We prove by contradiction. Assume there exists $\mathsf{Rev}(S_{i+1})\leq a<b<c\leq p^{*}$ satisfying $c=b+t\cdot(b-a)$ and $R(c)>R(b)+t\cdot(R(b)-R(a))$. Define $q(x):=1-F(x)$ and $\overline{q}:=q(b)+t\cdot\big{(}q(b)-q(a)\big{)}$. We first show that $\overline{q}\geq q(c)$: Since the revenue function is concave in the quantile space, hence, for all $q^{\prime}\geq\overline{q}$,
$R(q^{-1}(q^{\prime}))~{}\leq~{}R(b)+\frac{q(b)-q^{\prime}}{q(a)-q(b)}\cdot(R(b)-R(a))~{}\leq~{}R(b)+t\cdot(R(b)-R(a))~{}<~{}R(c).$
For every $q^{\prime}\geq\bar{q}$, we have $q^{\prime}\neq q(c)$. So. we must have $q(c)<\bar{q}$. This gives the desired contradiction:
$\frac{R(c)}{P_{i}}~{}>~{}$ $\frac{R(b)}{P_{i}}+\frac{t\cdot(R(b)-R(a))}{P_{i}}$
[…]
$-t(t+1)(b-a)(q(a)-q(b))$
[…]
The claim shows that $R_{i}(p)$ is half-concave, and therefore learnable in $[\mathsf{Rev}(S_{i+1}),1]$. However, in $[0,\mathsf{Rev}(S_{i+1})]$ the function is increasing but not necessarily concave. The sub-algorithm $\mathsf{SubAlg}$ will involve different cases to handle non-concavity inside interval $[0,\mathsf{Rev}(S_{i+1})]$. Depending on the value of $P_{i}$ and $F_{i}(r^{s}_{i})$, we run different algorithms. See Algorithm 7 for the pseudo-code of $\mathsf{SubAlg}$.
Input: Hidden function $R_{i}(p)$, interval $[\ell_{i},r_{i}]$, error parameter $\delta$
1 Fix $p_{j}=r^{s}_{j}$ for $j<i$ and $p_{j}=\hat{p}_{j}$ for $j>i$.
2 Test $N=C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=r^{s}_{i}$ to estimate $\hat{P}_{i}=\frac{\sum_{i\in[N]}\mathbf{1}[\text{Item not sold before }i]}{N}$.
3 if $\hat{P}_{i}<\frac{3}{4}\delta$ then
       Output $\hat{p}_{i}=\ell_{i}$ and Return. // Case 1: $P_{i}$ too small, any price in $[\ell_{i},r^{s}_{i}]$ works.
4
5
6Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=r^{s}_{i}$ to estimate $\hat{F}_{i}(r^{s}_{i}):=\frac{\sum\mathbf{1}[\text{Item not sold before }i+1]}{\sum\mathbf{1}[\text{Item not sold before }i]}$.
7 if $\hat{F}_{i}(r^{s}_{i})\geq 0.4$ then
       // Case 2: $F_{i}(r^{s}_{i})\geq 0.3$, estimate $\mathsf{Rev}(S_{i+1})$ accurately
8       Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=r^{s}_{i}$ to get $\widehat{\mathsf{Rev}}(S_{i+1}):=\frac{\sum\text{Revenue if item not sold before }i+1}{\sum\mathbf{1}[\text{Item not sold before }i+1]}$.
9       Call Algorithm 2 with input $R_{i}(p)$, interval $[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]$, $\delta$, and receive output $\hat{p}^{(1)}_{i}$.
10       Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=\hat{p}^{(1)}_{i}$. Let $\hat{R}_{i}(\hat{p}^{(1)}_{i})$ be the average reward.
11       Let $\hat{p}^{(2)}_{i}=\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})-\frac{\delta}{\hat{P}_{i}}\}$.
12       Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=\hat{p}^{(2)}_{i}$. Let $\hat{R}_{i}(\hat{p}^{(2)}_{i})$ be the average reward.
13       Let $\hat{p}_{i}=\arg\max_{p\in\{\hat{p}^{(1)}_{i},\hat{p}^{(2)}_{i}\}}\hat{R}_{i}(p)$.
14else
       // Case 3: $F_{i}(r^{s}_{i})\leq 0.5$, either $R_{i}(p)$ is half 2-concave or $R_{i}(r^{s}_{i})$ is good.
15       Call $\mathsf{Alg}$ and receive output $\hat{p}^{(3)}_{i}$.
16       Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=\hat{p}^{(3)}_{i}$. Let $\hat{R}_{i}(\hat{p}^{(3)}_{i})$ be the average reward.
17       Let $\hat{p}^{(4)}_{i}=r^{s}_{i}$.
18       Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=\hat{p}^{(4)}_{i}$. Let $\hat{R}_{i}(\hat{p}^{(4)}_{i})$ be the average reward.
19       Let $\hat{p}_{i}=\arg\max_{p\in\{\hat{p}^{(3)}_{i},\hat{p}^{(4)}_{i}\}}\hat{R}_{i}(p)$.
Output: $\hat{p}_{i}$
ALGORITHM 7 Sub-algorithm $\mathsf{SubAlg}$
There are three cases in Algorithm 7. We first briefly discuss the intuition behind each case.
Case 1 ($P_{i}$ is small): Since the difference between $R_{i}(\hat{p}^{*}_{i})$ and any other $R_{i}(p)$ is upper bounded by $P_{i}$, playing any price in $[\ell_{i},r^{s}_{i}]$ does not incur a big regret. Therefore, taking $\hat{p}_{i}$ to be an arbitrary price in $[\ell_{i},r^{s}_{i}]$ is feasible.
Case 2 ($F_{i}(r^{s}_{i})$ is large): In this case, the value of $\mathsf{Rev}(S_{i+1})$ can be estimated accurately. Then, we divide interval $[\ell_{i},r^{s}_{i}]$ from the upper bound of $\mathsf{Rev}(S_{i+1})$ into two parts. For the part on the right, 3.9 guarantees that $R_{i}(p)$ is half-concave, and we call the single-buyer algorithm to find a candidate price. For the part on the left, the constrained Lipschitzness from 3.9 suggests that we can mark the lower bound of $\mathsf{Rev}(S_{i+1})$ as a candidate price. Finally, we take a better one from two candidate prices to be $\hat{p}_{i}$.
Case 3 ($F_{i}(r^{s}_{i})$ is small): In this case, we find two candidate prices according to the value of $r^{s}_{i}$ and $\mathsf{Rev}(S_{i+1})$. If $r^{s}_{i}\geq\mathsf{Rev}(S_{i+1})$, the following lemma gives the first candidate price:
Lemma 3.10.
If $F_{i}(r^{s}_{i})\leq 0.5$ and $r^{s}_{i}\geq\mathsf{Rev}(S_{i+1})$, there exists an algorithm $\mathsf{Alg}$ that runs $O(\delta^{-2}\log^{2}T)$ rounds with price vectors $(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i}\in[\ell_{i},r^{s}_{i}],\hat{p}_{i+1},\cdots,\hat{p}_{n})$ and outputs $\hat{p}_{i}$ that satisfies with probability $1-T^{-6}$ that $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$.
3.10 suggests that $R_{i}(p)$ is learnable when $F_{i}(r^{s}_{i})$ is small and $r^{s}_{i}\geq\mathsf{Rev}(S_{i+1})$ holds. The proof of 3.10 requires ideas of generalized half-concavity. For consistency, we first assume the correctness of 3.10, and defer the details to Section 3.8.
On the other hand, if instead $\mathsf{Rev}(S_{i+1})>r^{s}_{i}$ holds, we use $r^{s}_{i}$ to be the second candidate price. The main reason is from the following claim:
Claim 3.11.
Assume $p^{*}_{i}\in[\ell_{i},r_{i}]$, then there must be $\hat{p}^{*}_{i}\geq\mathsf{Rev}(S_{i+1})$.
Proof.
Suppose $\hat{p}^{*}_{i}<\mathsf{Rev}(S_{i+1})$. Then, we have $p^{*}_{i}\geq\mathsf{Opt}(S_{i+1})\geq\mathsf{Rev}(S_{i+1})>\hat{p}^{*}_{i}$, where the first inequality is implied by the observation that
$R^{*}_{i}(p^{*}_{i})-R^{*}_{i}(\mathsf{Opt}(S_{i+1}))~{}=~{}P_{i}\cdot(1-F_{i}(p^{*}_{i}))\cdot(p^{*}_{i}-\mathsf{Opt}(S_{i+1}))~{}\geq~{}0.$
Since we have $\hat{p}^{*}_{i},p^{*}_{i}\in[\ell_{i},r_{i}]$, there must be $\mathsf{Rev}(S_{i+1})\in[\ell_{i},r_{i}]$. This leads to a contradiction: 3.9 guarantees $R_{i}(p)$ is increasing in $[0,\mathsf{Rev}(S_{i+1})]$, which is in contrast to the assumption that $\hat{p}^{*}_{i}<\mathsf{Rev}(S_{i+1})$. Therefore, there must be $\hat{p}^{*}_{i}\geq\mathsf{Rev}(S_{i+1})$. ∎
3.11 together with the inequality $\mathsf{Rev}(S_{i+1})>r^{s}_{i}$ implies $\hat{p}^{*}_{i}\in[r^{s}_{i},r_{i}]$. Now the constrained Lipschitzness of $R_{i}(p)$ implies $R_{i}(\hat{p}^{*}_{i})-R_{i}(r^{s}_{i})\leq\hat{p}^{*}_{i}-r^{s}_{i}\leq{T^{-100}}$. Therefore, it’s sufficient to take $r^{s}_{i}$ as the second candidate price. Then, the algorithm takes better of the two prices for Case 3.
Now, we prove 3.5. Our proof starts from the following concentration bounds for $\mathsf{SubAlg}$:
Claim 3.12.
In Algorithm 7, with probability $1-T^{-6}$ the following bounds simultaneously hold:
  • $|\hat{P}_{i}-P_{i}|\leq\frac{1}{4}\delta$.
  • $|\hat{F}_{i}(r^{s}_{i})-F_{i}(r^{s}_{i})|\leq 0.1$.
  • $|\widehat{\mathsf{Rev}}(S_{i+1})-\mathsf{Rev}(S_{i+1})|\leq\frac{\delta}{\hat{P}_{i}}$.
  • $|\hat{R}_{i}(\hat{p}^{(j)}_{i})-R_{i}(\hat{p}^{(j)}_{i})|\leq\frac{1}{2}\delta$ for all $j\in\{1,2,3,4\}$.
Proof.
We show each statement holds with probability $1-T^{-7}$ and the union bound guarantees all statements hold with probability $1-T^{-6}$.
(i) $|\hat{P}_{i}-P_{i}|\leq\frac{1}{4}\delta$: $\hat{P}_{i}$ is estimated with $N=C\cdot\delta^{-2}\log T$ samples. By Hoeffding’s Inequality, we have
$\mbox{\rm\bf Pr}\left[|\hat{P}_{i}-P_{i}|>\frac{1}{4}\delta\right]~{}\leq~{}2\exp(-2N\cdot\frac{\delta^{2}}{16})~{}=~{}2T^{C/8}~{}\leq T^{-7},$
where the last inequality holds when $C$ is a constant greater than $60$.
(ii) $|\hat{F}_{i}(r^{s}_{i})-F_{i}(r^{s}_{i})|\leq 0.1$. We assume $|\hat{P}_{i}-P_{i}|\leq\frac{1}{4}\delta$ holds. Then, the algorithm estimates $\hat{F}_{i}(r^{s}_{i})$ only when $P_{i}\geq\hat{P}_{i}-\frac{1}{4}\delta\geq\frac{1}{2}\delta$.
In the algorithm, we run $N^{\prime}=C\frac{\log T}{\delta^{2}}$ rounds to estimate $\hat{F}_{i}(r^{s}_{i})$. However, a single round becomes a valid sample only when the item is unsold until buyer $i$. We claim that with probability $1-T^{-7}$ we have at least $N=400\log T$ samples. This is true because by Hoeffding’s Inequality, we have
$\mbox{\rm\bf Pr}\left[N<400\log T\right]~{}<$ $~{}\mbox{\rm\bf Pr}\left[\left|\frac{N}{N^{\prime}}-P_{i}\right|>P_{i}-\frac{400\log T}{N^{\prime}}\right]$
[…]
Here, the second inequality is from $P_{i}\geq\frac{1}{2}\delta$ and $\frac{400\log T}{N^{\prime}}=\frac{400\delta^{2}}{C}<\frac{1}{4}\delta$ when $C$ is a constant greater than $1600$, and the last inequality also holds when $C>1600$.
Next, we show $|\hat{F}_{i}(r^{s}_{i})-F_{i}(r^{s}_{i})|\leq 0.1$ assuming we have at least $N=400\log T$ samples. By Hoeffding’s Inequality, we have
$\mbox{\rm\bf Pr}\left[|\hat{F}_{i}(r^{s}_{i})-F_{i}(r^{s}_{i})|>0.1\right]~{}\leq~{}2\exp(-2N\cdot 0.01)~{}=~{}2T^{-8}~{}<~{}T^{-7}.$
(iii) $|\widetilde{\mathsf{Rev}}(S_{i+1})-\mathsf{Rev}(S_{i+1})|\leq\frac{\delta}{\hat{P}_{i}}$. We assume $|\hat{P}_{i}-P_{i}|\leq\frac{1}{4}\delta$ and $|\hat{F}_{i}(r^{s}_{i})-F_{i}(r^{s}_{i})|\leq 0.1$ holds. The algorithm estimates $\widetilde{\mathsf{Rev}}(S_{i+1})$ only when $\hat{P}_{i}\geq\frac{3}{4}\delta$ and $\hat{F}_{i}(r^{s}_{i})\geq 0.4$ hold. In this case, $|\hat{P}_{i}-P_{i}|\leq\frac{1}{4}\delta$ and $\hat{P}_{i}\geq\frac{3}{4}\delta$ give $\hat{P}_{i}\in[\frac{3}{4}P_{i},\frac{3}{2}P_{i}]$. Besides, we have $P_{i}\geq\hat{P}_{i}-\frac{1}{4}\delta\geq\frac{1}{2}\delta$ and $F_{i}(r^{s}_{i})\geq\hat{F}_{i}(r^{s}_{i})-0.1\geq 0.3$.
In the algorithm, we run $N^{\prime}=C\frac{\log T}{\delta^{2}}$ rounds to estimate $\widetilde{\mathsf{Rev}}(S_{i+1})$. However, a single round becomes a valid sample only when the item is unsold until buyer $i+1$. We claim that with probability $1-T^{-7}$ we have at least $N=\frac{9\log TP^{2}_{i}}{\delta^{2}}$ samples. This is true because by Hoeffding’s Inequality, we have
$\mbox{\rm\bf Pr}\left[N<\frac{9\log TP^{2}_{i}}{\delta^{2}}\right]~{}\leq$ $~{}\mbox{\rm\bf Pr}\left[\left|\frac{N}{N^{\prime}}-P_{i+1}\right|>P_{i+1}-\frac{9\log TP^{2}_{i}}{N^{\prime}\cdot\delta^{2}}\right]$
[…]
Here, the second inequality is from $P_{i+1}=P_{i}F_{i}(r^{s}_{i})\geq 0.3P_{i}$, and $\frac{9\log TP_{i}^{2}}{N^{\prime}\delta^{2}}=\frac{9P_{i}^{2}}{C}<0.1P_{i}$, when $C$ is a constant greater than $90$. The second to the last inequality uses the fact that $P_{i}\geq\frac{1}{2}\delta$, and the last inequality is true when $C$ is a constant greater than $800$.
Next, we show $|\widetilde{\mathsf{Rev}}(S_{i+1})-\mathsf{Rev}(S_{i+1})|\leq\frac{\delta}{\hat{P}_{i}}$ assuming we have at least $N=400\log T$ samples. By Hoeffding’s Inequality, we have
$\mbox{\rm\bf Pr}\left[|\widetilde{\mathsf{Rev}}(S_{i+1})-\mathsf{Rev}(S_{i+1})|>\frac{\delta}{\hat{P}_{i}}\right]~{}\leq$ $~{}\mbox{\rm\bf Pr}\left[|\widetilde{\mathsf{Rev}}(S_{i+1})-\mathsf{Rev}(S_{i+1})|>\frac{\delta}{1.5P_{i}}\right]$
[…]
(iii) $|\hat{R}_{i}(\hat{p}^{(j)}_{i})-R_{i}(\hat{p}^{(j)}_{i})|\leq\frac{1}{2}\delta$ for $j\in\{1,2,3,4\}$. We only prove the inequality when $j=1$ and the remaining cases are identical. $\hat{R}_{i}(\hat{p}^{(1)}_{i})$ is estimated via $N=C\cdot\frac{\log T}{\delta^{2}}$ samples. By Hoeffding’s Inequality, we have
$\mbox{\rm\bf Pr}\left[|\hat{R}_{i}(\hat{p}^{(j)}_{i})-R_{i}(\hat{p}^{(j)}_{i})|>\frac{1}{2}\delta\right]~{}\leq~{}2\exp(-2N\cdot 0.25\delta^{2})~{}=~{}2T^{-0.5C}~{}<~{}T^{-7},$
where the last inequality holds when $C$ is a constant greater than $16$. ∎
With the help of 3.12, we can give the formal proof of 3.5:
Proof of 3.5.
The bound of number of tested rounds in and the constraint that $p_{i}\in[\ell_{i},r^{s}_{i}]$ are directly guaranteed by Algorithm 7. It remains to show $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$. We prove the inequality separately for the following cases:
Case 1: $\hat{P}_{i}<\frac{3}{4}\delta$ (Line 3-4). In this case, $\hat{P}_{i}<\frac{3}{4}\delta$ implies $P_{i}\leq\delta$, and $R_{i}(\hat{p}^{*}_{i})-R_{i}(p)\leq P_{i}\leq\delta$ holds for any price $p$, so taking $\hat{p}_{i}=\ell_{i}$ is feasible.
Case 2: $\hat{P}_{i}\geq\frac{3}{4}\delta$ while $\hat{F}_{i}(r^{s}_{i})\geq 0.4$ (Line 7-12). Consider the value of $\hat{p}^{*}_{i}$. If $\hat{p}^{*}_{i}\in[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]$, there must be $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(1)}_{i})\leq\delta$. This is true because 3.9 guarantees $[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]\subseteq[\mathsf{Rev}(S_{i+1}),1]$ is half-concave, and 2.7 guarantees that the single-buyer algorithm outputs $\hat{p}^{(1)}_{i}$ that satisfies $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(1)}_{i})\leq\delta$ with probability $1-T^{-6}$.
If $\hat{p}^{*}_{i}\notin[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]$, there must be $\hat{p}^{*}_{i}\in[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})-\frac{\delta}{\hat{P}_{i}}\},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\}]$. This is true because simultaneously we have $\hat{p}^{*}_{i}\geq\ell_{i}$ from definition of $\hat{p}^{*}_{i}$, $\hat{p}^{*}_{i}\geq\mathsf{Rev}(S_{i+1})\geq\widehat{\mathsf{Rev}}(S_{i+1})-\frac{\delta}{\hat{P}_{i}}$ from 3.11 and 3.12, and $\hat{p}^{*}_{i}\leq\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}$ from the assumption $\hat{p}^{*}_{i}\notin[\max\{\ell_{i},\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}}\},r_{i}]$. Then, the constrained Lipschitzness from 3.9 gives
$R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(2)}_{i})\leq P_{i}(\hat{p}^{*}_{i}-\hat{p}^{(2)}_{i})\leq P_{i}\big{(}(\widehat{\mathsf{Rev}}(S_{i+1})+\frac{\delta}{\hat{P}_{i}})-(\widehat{\mathsf{Rev}}(S_{i+1})-\frac{\delta}{\hat{P}_{i}})\big{)}\leq\frac{2P_{i}\delta}{\hat{P}_{i}}\leq 3\delta,$
where the last inequality holds because $\hat{P}_{i}\geq\frac{3}{4}\delta$ and $|\hat{P}_{i}-P_{i}|\leq\frac{1}{4}\delta$ imply $\frac{P_{i}}{\hat{P}_{i}}\leq\frac{4}{3}<1.5$.
Therefore, at least one of the inequalities $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(1)}_{i})\leq 3\delta$ or $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(2)}_{i})\leq 3\delta$ holds. If $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(1)}_{i})\leq 3\delta$ holds, we have
$R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})~{}\leq~{}R_{i}(\hat{p}^{*}_{i})-\hat{R}_{i}(\hat{p}_{i})+\frac{\delta}{2}~{}\leq~{}R_{i}(\hat{p}^{*}_{i})-\hat{R}_{i}(\hat{p}^{(1)}_{i})+\frac{\delta}{2}~{}\leq~{}R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(1)}_{i})+\delta~{}\leq~{}4\delta.$
Symmetrically, if $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(2)}_{i})\leq 3\delta$, the same argument still gives $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$. Therefore, $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ holds in Case 2.
Cases 3: $\hat{P}_{i}\geq\frac{3}{4}\delta$ while $\hat{F}_{i}(r^{s}_{i})<0.4$ (Line 14-18). This is sufficient to show $F_{i}(r^{s}_{i})<0.5$.
In this case, the algorithm prepares two candidate prices. If $\mathsf{Rev}(S_{i+1})\leq r^{s}_{i}$, 3.10 guarantees that with probability $1-T^{-6}$ we get $\hat{p}^{(3)}_{i}$ such that $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(3)}_{i})\leq\delta$. Otherwise, there must be $\mathsf{Rev}(S_{i+1})>r^{s}_{i}$. In this case, 3.11 guarantees $r^{s}_{i}<\mathsf{Rev}(S_{i+1})\leq\hat{p}^{*}_{i}$, and the constrained Lipschitzness of $R_{i}(p)$ guarantees
$R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(4)}_{i})~{}\leq~{}\hat{p}^{*}_{i}-\hat{p}^{(4)}_{i}~{}=~{}\hat{p}^{*}_{i}-r^{s}_{i}~{}\leq~{}{T^{-100}}~{}<~{}\delta.$
Therefore, at least one of the inequalities $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(3)}_{i})\leq\delta$ or $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}^{(4)}_{i})\leq\delta$ holds. An argument similar to Case 1 is sufficient to show $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$.
To conclude, $R_{i}(\hat{p}^{*}_{i})-R_{i}(\hat{p}_{i})\leq 4\delta$ holds in all cases. The final step is to check the success probability of our proof. In this proof, we require the correctness of 3.12, Algorithm 2 in Case 2, and $\mathsf{Alg}$ in Case 3. Each of them fails with probability $T^{-6}$. By the union bound, the whole proof succeeds with probability $1-3T^{-6}>1-T^{-5}$. ∎
3.8 3.10 via Generalized Half-Concavity
In this subsection, we prove the missing 3.10, which is restated for convenience.
See 3.10
Our algorithm begins from the following observation: Recall that $R_{i}(p)=\mathsf{Pre}(S_{i})+P_{i}(p\cdot(1-F_{i}(p))+\mathsf{Rev}(S_{i+1})\cdot F_{i}(p))$. Then,
$R^{\prime}_{i}(p)~{}=~{}P_{i}\big{(}1-F_{i}(p)+(\mathsf{Rev}(S_{i+1})-p)\cdot f_{i}(p)\big{)}.$
When $F_{i}(r^{s}_{i})\leq 0.5$, we have $R^{\prime}_{i}(p)\geq P_{i}(1-F_{i}(p))\geq P_{i}(1-F_{i}(r^{s}_{i}))\geq\frac{1}{2}P_{i}$ for $p\in[0,\mathsf{Rev}(S_{i+1})]$, while $R^{\prime}_{i}(p)\leq P_{i}(1-F_{i}(p))\leq P_{i}$ for $p\in[\mathsf{Rev}(S_{i+1}),1]$. This observation shows $R_{i}(p)$ is growing sufficiently fast in $[0,\mathsf{Rev}(S_{i+1})]$ as compared to $[\mathsf{Rev}(S_{i+1}),1]$, leading to some kinds of “concavity” property in $[0,\mathsf{Rev}(S_{i+1})]$. Inspired by this observation, we give the following definition of the generalized half-concavity and show how it relates to $R_{i}(p)$:
Definition 3.13.
[Half $\lambda$-Concavity] Given function $R(p)$ and interval $[\ell,r]$. For $\lambda\geq 1$, we say $R(p)$ is half $\lambda$-concave on $[\ell,r]$ if the following conditions hold:
  • $R(p)$ is single-peaked in $[\ell,r]$.
  • Let $p^{*}=\arg\!\max_{[\ell,r]}R(p)$. For $p\in[\ell,p^{*}]$, $R(p^{*})-R(p)\leq p^{*}-p$.
  • For $a,b\in[\ell,p^{*})$ satisfying $a<b$, we have
    $\frac{R(b)-R(a)}{b-a}\geq\frac{1}{\lambda}\cdot\frac{R(p^{*})-R(b)}{p^{*}-b}.$
The intuition of defining half $\lambda$-concavity is to capture the case when $R(p)$ is concave, or the case when $R(p)$ has slope between $[k,\lambda k]$, or a mixture of these two. The following lemma suggests $R_{i}(p)$ is captured by this half $\lambda$-concavity when $F_{i}(r^{s}_{i})$ is not too large.
Lemma 3.14.
If $F_{i}(r^{s}_{i})<\frac{1}{2}$ and $r^{s}_{i}>\mathsf{Rev}(S_{i+1})$, $R_{i}(p)$ is half 2-concave in $[\ell_{i},r_{i}]$.
Proof.
We prove three properties one by one.
(i) Single-peakedness: This is guaranteed by 3.9.
(ii) Lipschitzness with respect to $p^{*}$: 3.11 gives $\hat{p}^{*}_{i}\geq\mathsf{Rev}(S_{i+1})$. Then, the constrained Lipschitzness from 3.9 is sufficient to give the inequality $R_{i}(\hat{p}^{*}_{i})-R_{i}(p)\leq\hat{p}^{*}_{i}-p$ for all $p\in[\ell_{i},\hat{p}^{*}_{i}]$.
(iii) Weakened Concavity: The proof is based on the following three facts:
  • Fact 1: For $x<y\leq\mathsf{Rev}(S_{i+1})$, we have
[…]
  • Fact 2: For $x\leq\mathsf{Rev}(S_{i+1})$ and $y>\mathsf{Rev}(S_{i+1})$, the constrained Lipschitzness in 3.9 gives
    $\frac{R_{i}(y)-R_{i}(x)}{y-x}~{}\leq~{}P_{i}.$
  • Fact 3: For $x,y\in[\ell_{i},r_{i}]$ that satisfies $\mathsf{Rev}(S_{i+1})\leq x<y<\hat{p}^{*}_{i}$, we claim that
    $\frac{R_{i}(y)-R_{i}(x)}{y-x}~{}\geq~{}\frac{R_{i}(\hat{p}^{*}_{i})-R_{i}(y)}{\hat{p}^{*}_{i}-y}.$
    If a feasible pair of $(x,y)$ that satisfies $x,y\in[\ell_{i},r_{i}]$ and $\mathsf{Rev}(S_{i+1})\leq x<y<\hat{p}^{*}_{i}$ exists, there must be $\hat{p}^{*}_{i}>\ell_{i}$, which implies that $R_{i}(p)$ is not decreasing in $[\ell_{i},r_{i}]$, and there must be $\mathsf{Rev}(S_{i+1})\leq\hat{p}^{*}_{i}\leq\arg\max_{p\in[0,1]}R_{i}(p)$. Then, the half-concavity of $R_{i}(p)$ in $[\mathsf{Rev}(S_{i+1}),1]$ guarantees that $R_{i}(p)$ is concave in $[\mathsf{Rev}(S_{i+1}),\hat{p}^{*}_{i}]$. This gives the inequality in Fact 3.
Now, we divide the proof into the following three cases according to the value of $a$ and $b$:
  • Case 1: $a<b\leq\mathsf{Rev}(S_{i+1})$. In this case, we have
    $\frac{R_{i}(b)-R_{i}(a)}{b-a}~{}\geq~{}\frac{1}{2}P_{i}~{}\geq~{}\frac{1}{2}\cdot\frac{R_{i}(\hat{p}^{*}_{i})-R_{i}(b)}{\hat{P}^{*}_{i}-b},$
    where the first inequality is from Fact 1 and the second inequality is from Fact 2.
  • Case 2: $a<\mathsf{Rev}(S_{i+1})$, while $b>\mathsf{Rev}(S_{i+1})$. In this case, Fact 1 and 2 gives
    $\frac{R_{i}(\mathsf{Rev}(S_{i+1}))-R_{i}(a)}{\mathsf{Rev}(S_{i+1})-a}~{}\geq~{}\frac{1}{2}P_{i}~{}\geq~{}\frac{1}{2}\cdot\frac{R_{i}(\hat{p}^{*}_{i})-R_{i}(b)}{\hat{p}^{*}_{i}-b}.$
    On the other hand, Fact 3 gives
    $\frac{R_{i}(b)-R_{i}(\mathsf{Rev}(S_{i+1}))}{b-\mathsf{Rev}(S_{i+1})}~{}\geq~{}\frac{R_{i}(\hat{p}^{*}_{i})-R_{i}(b)}{\hat{p}^{*}_{i}-b}~{}\geq~{}\frac{1}{2}\cdot\frac{R_{i}(\hat{p}^{*}_{i})-R_{i}(b)}{\hat{p}^{*}_{i}-b}.$
    Therefore,
    $\frac{R_{i}(b)-R_{i}(a)}{b-a}\geq\min\left\{\frac{R_{i}(\mathsf{Rev}(S_{i+1}))-R_{i}(a)}{\mathsf{Rev}(S_{i+1})-a},\frac{R_{i}(b)-R_{i}(\mathsf{Rev}(S_{i+1}))}{b-\mathsf{Rev}(S_{i+1})}\right\}\geq\frac{1}{2}\cdot\frac{R_{i}(\hat{p}^{*}_{i})-R_{i}(b)}{\hat{p}^{*}_{i}-b}.$
  • Case 3: $\mathsf{Rev}(S_{i+1})\leq a<b<\hat{p}^{*}_{i}$. In this case, Fact 3 gives
    $\frac{R_{i}(b)-R_{i}(a)}{b-a}~{}\geq~{}\frac{R_{i}(\hat{p}^{*}_{i})-R_{i}(b)}{\hat{p}^{*}_{i}-b}~{}\geq~{}\frac{1}{2}\cdot\frac{R_{i}(\hat{p}^{*}_{i})-R_{i}(b)}{\hat{p}^{*}_{i}-b}.\qed$
3.14 shows that $R_{i}(p)$ is generalized half-concave. Then, the following lemma suggests that the function is learnable with the help of half $\lambda$-concavity:
Lemma 3.15.
If $R_{i}(p)$ is half $\lambda$-concave, then there exists an algorithm that takes parameter $\lambda$, interval $[\ell,r]$, error $\epsilon\geq\frac{1}{\sqrt{T}}$ as input, tests $O(\frac{\lambda^{2}\log^{2}T}{\epsilon^{2}})$ rounds with price vectors $(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i}\in[\ell,r-{T^{-100}}],\hat{p}_{i+1},\cdots,\hat{p}_{n})$, and outputs $\hat{p}_{i}$ that satisfies with probability $1-T^{-6}$ that $\max_{p\in[\ell,r]}R_{i}(p)-R_{i}(\hat{p}_{i})\leq\epsilon$.
3.15 is a generalization of Algorithm 2 for a single buyer. Observe that Algorithm 2 only needs a weaker inequality $\frac{R(b)-R(a)}{b-a}\geq\frac{R(p^{*})-R(b)}{p^{*}-b}$. Then, generalizing Algorithm 2 by taking $\lambda$ into consideration gives the desired algorithm for 3.15.
Observe that combining 3.15 with 3.14 directly proves 3.10. It only remains to prove 3.15.
Input: Hidden Function $R_{i}(p)$, Interval $[\ell,r]$, Concavity Parameter $\lambda$, Error Parameter $\epsilon$
1 Fix $p_{j}=r^{s}_{j}$ for $j<i$ and $p_{j}=\hat{p}_{j}$ for $j>i$.
2 Let $\delta=\frac{\epsilon}{100\lambda}$ be the scaled error parameter, $\ell_{1}\leftarrow\ell,r_{1}\leftarrow r$, $j=1$.
3 while $r_{j}-\ell_{j}>\delta$ do
4       Let $a_{j}\leftarrow\frac{2\ell_{j}+r_{j}}{3},b_{j}\leftarrow\frac{\ell_{j}+2r_{j}}{3}$.
5       Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p_{i}=a_{j}$. Let $\hat{R}_{i}(a_{j})$ be the average reward.
6       Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p_{i}=b_{j}$. Let $\hat{R}_{i}(b_{j})$ be the average reward.
7       if $\hat{R}_{i}(a_{j})<\hat{R}_{i}(b_{j})-2\delta$ then
8            Let $\ell_{j+1}\leftarrow a_{j},r_{j+1}\leftarrow r_{j}$
9      else
10            Let $\ell_{j+1}\leftarrow\ell_{j},r_{j+1}\leftarrow b_{j}$
11      $j\leftarrow j+1$
12Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with price $p_{i}=\ell_{j}$. Let $\hat{R}_{i}(\ell_{j})$ be the average reward.
13 Let $\hat{p}_{i}=\arg\max_{p\in P}\hat{R}_{i}(p)$, where $P=\{p:p\text{ is tested}\}$.
Output: $\hat{p}_{i}$.
ALGORITHM 8 Generalized Algorithm for Half $\lambda$-Concave Functions
Proof of 3.15.
Algorithm 8 is the desired algorithm for 3.15. The algorithm is almost the same as Algorithm 2. The only difference is that extra care is needed for the extra parameter $\lambda$.
We first show that Algorithm 8 is feasible: it tests $O(\frac{\lambda^{2}\log^{2}T}{\epsilon^{2}})$ rounds with $p_{i}\in[\ell,r-{T^{-100}}]$.
The main body of Algorithm 8 is a while loop that maintains an interval to be tested. Assume the while loop stops when $j=k+1$. Then, for $j\in[k]$, the length of $[\ell_{j+1},r_{j+1}]$ is two-thirds of $[\ell_{j},r_{j}]$, and we have $r_{k}-\ell_{k}>\delta$. Therefore, $k=O(\log\frac{1}{\delta})=O(\log T)$. For interval $[\ell_{j},r_{j}]$, two prices $a_{j},b_{j}$ are tested for $O(\frac{\log T}{\delta^{2}})$ times, so in total Algorithm 8 tests $O(\frac{\lambda^{2}\log^{2}T}{\epsilon^{2}})$ rounds. To show every tested price is in $[\ell,r-{T^{-100}}]$, let $p_{i}$ be a price tested in Algorithm 8. $p_{i}\in[\ell,r]$ directly follows from the algorithm. For the inequality $p_{i}\leq r-{T^{-100}}$, observe that the largest tested price must be $p_{i}=b_{j}$ for some $j\in[k]$, and $b_{j}\leq r-{T^{-100}}$ is guaranteed by the fact that $r_{j}-b_{j}\geq\frac{1}{3}\delta\gg{T^{-100}}$.
It only remains to show $\max_{p\in[\ell,r]}-R_{i}(\hat{p}_{i})\leq\epsilon$. Define $p^{*}_{\ell,r}:=\arg\max_{p\in[\ell,r]}R_{i}(p)$ to be the optimal price in $[\ell,r]$. We will show that $R_{i}(p^{*}_{\ell,r})-R_{i}(\hat{p}_{i})\leq 6\lambda\delta<\epsilon$ holds with probability $1-T^{-6}$. We start the proof by introducing the following claim:
Claim 3.16.
In Algorithm 8, $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ simultaneously holds for every tested price $p$ with probability at least $1-T^{-6}$.
Proof.
For a single tested price $p$, $\hat{R}_{i}(p)$ is estimated by calculating the average of $N=C\cdot\frac{\log T}{\delta^{2}}$ samples. By Hoeffding’s Inequality,
$\mbox{\rm\bf Pr}\left[|\hat{R}_{i}(p)-R_{i}(p)|>\delta\right]~{}\leq~{}2\exp\left(-2N\delta^{2}\right)~{}=~{}2T^{-2C}~{}<~{}T^{-7}.$
The last inequality holds when $C$ is a constant greater than $4$. Then, we have $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ holds with probability $1-T^{-7}$ for a single tested price $p$.
Notice that Algorithm 8 can’t test more than $T$ prices. By the union bound, $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ simultaneously holds for all tested prices with probability $1-T^{-6}$. ∎
Now, we prove $R_{i}(p^{*}_{\ell,r})-R_{i}(\hat{p}_{i})\leq 6\lambda\delta$ assuming 3.16 holds. We demonstrate the proof by discussing the following two cases. The first case is that there exists $j\in[k]$, such that $p^{*}_{\ell,r}$ falls in $[\ell_{j},r_{j}]$ but not in $[\ell_{j+1},r_{j+1}]$. In this case, there must be $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$ and $[b_{j},r_{j}]$ is dropped. If not, we have $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$ and $[\ell_{j},a_{j}]$ is dropped. However, the inequality $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$ together with 3.16 gives $R(a_{j})<R(b_{j})$. Then, the single-peakedness of $R_{i}(p)$ gives $p^{*}_{\ell,r}\geq a_{j}$, which is in contrast to the assumption that $p^{*}_{\ell,r}\notin[\ell_{j+1},r_{j+1}]$. Therefore, we have $\hat{R}_{i}(a_{j})\geq\hat{R}_{i}(b_{j})-2\delta$, and the assumption $p^{*}_{\ell,r}\notin[\ell_{j+1},r_{j+1}]$ implies $p^{*}_{\ell,r}\in[b_{j},r_{j}]$. Then, the half $\lambda$-concavity of $R_{i}(p)$ in $[a_{j},p^{*}_{\ell,r}]$ guarantees
$R_{i}(p^{*}_{\ell,r})-R_{i}(b_{j})~{}\leq~{}\frac{R_{i}(b_{j})-R_{i}(a_{j})}{b_{j}-a_{j}}\cdot\lambda(\hat{p}^{*}_{\ell,r}-b_{j})~{}\leq~{}\lambda\big{(}(\hat{R}_{i}(b_{j})+\delta)-(\hat{R}_{i}(a_{j})-\delta)\big{)}~{}\leq~{}4\lambda\delta.$
The second case is that the desired $j$ in the first case doesn’t exist. In this case, the only possibility is that $p^{*}_{\ell,r}\in[\ell_{k+1},r_{k+1}]$. Then, the Lipschitzness of function $R_{i}(p)$ w.r.t. $p^{*}_{\ell,r}$ guarantees that $R_{i}(p^{*}_{\ell,r})-R_{i}(\ell_{k})\leq\delta<4\lambda\delta$.
In both cases, we find a tested price $p^{\prime}$ satisfying $R_{i}(p^{\prime})\geq R_{i}(p^{*}_{\ell,r})-4\lambda\delta$. Then,
$R_{i}(\hat{p})~{}\geq~{}\hat{R}_{i}(\hat{p})-\delta~{}\geq\hat{R}_{i}(p^{\prime})-\delta~{}\geq~{}R_{i}(p^{\prime})-2\delta~{}\geq~{}R_{i}(p^{*}_{\ell,r})-2\delta-4\lambda\delta~{}\geq~{}6\lambda\delta.\qed$
4 $n$ Buyers with General Distributions
In this section, we provide the $\widetilde{O}(\mathsf{poly}(n)T^{\frac{2}{3}})$ regret algorithm for BSPP problem with general distributions. The setting in this section is the same as Section 3, but with no extra regularity assumption. Our result generalizes the $\widetilde{O}(T^{2/3})$ single-buyer result studied in [KL03].
Theorem 4.1.
For BSPP problem with $n$ buyers, there is an algorithm with $O(n^{5/3}T^{2/3}\log T)$ regret.
Algorithm Overview.
Comparing to the regular buyers setting, one extra challenge for general buyers setting is that the revenue function in this case is no longer single-peaked, leading to the failure of the idea of confidence intervals. Therefore, the first step of our algorithm is to discretize the value space: it can be shown that rounding buyers’ valuations to multiples of $\epsilon$ only brings an extra $\epsilon$ loss. The idea of discretization was first used in [KL03] for single buyer setting. We will show that the same idea also works for sequential buyers. In this section, we take the discretization parameter $\epsilon=n^{5/3}T^{-1/3}$, and the accumulated regret from the discretization step would be $T\cdot O(\epsilon)=O(n^{5/3}T^{2/3})$.
After discretizing the value space, it suffices to solve BSPP for general distributions with discrete support. The main idea of our algorithm is almost identical to the algorithm for regular buyers - we design a subroutine that shrinks the possible candidates of $p^{*}_{i}$ while keeping a low regret inside the new set of candidates. Calling this subroutine $O(\log T)$ times with a halving error parameter gives the desired algorithm. Comparing to the regular distributions setting, the idea of confidence intervals does not work for discrete distributions, because the candidates may be discontinuous. To fix this issue, we use $n$ price sets to maintain possible candidate prices.
The detailed steps of our algorithm also follows the ideas for regular distributions: we first find a group of approximately prices $\hat{p}_{1},\cdots,\hat{p}_{n}$, and then use these prices as a benchmark to update the candidate prices sets. Since the single-peakedness no longer holds for general distributions, instead of doing binary search, we simply enumerate all remaining prices to update the candidate sets.
4.1 BSPP is Discretizable
We start our proof by first showing BSPP is discretizable. Specifically, we have the following lemma:
Lemma 4.2.
Consider a sequential posted pricing problem. Let $k$ be an integer and $\epsilon=\frac{1}{k}$ be the discretization error. If we restrict the space of the prices to be $P=\{(p_{1},\cdots,p_{n}):p_{i}=j\cdot\epsilon,j\in\mathbb{Z}^{+}\}$, then
$\max_{(p_{1},\cdots,p_{n})\in[0,1]^{n}}R(p_{1},\cdots,p_{n})~{}-~{}\max_{(p_{1},\cdots,p_{n})\in P}R(p_{1},\cdots,p_{n})~{}\leq~{}\epsilon.$
Proof of 4.2.
Let $p^{*}_{1},\cdots,p^{*}_{n}$ be the optimal prices for $R(p_{1},\cdots,p_{n})$, and let $p^{\prime}_{i}=\lfloor\frac{p^{*}_{i}}{\epsilon}\rfloor\cdot\epsilon$. To prove 4.2, we only need to show
$R(p^{*}_{1},\cdots,p^{*}_{n})~{}-~{}R(p^{\prime}_{1},\cdots,p^{\prime}_{n})~{}\leq~{}\epsilon.$
Let $E_{i}$ be the event that an item is sold to buyer $i$ with price vector $(p^{\prime}_{1},\cdots,p^{\prime}_{n})$, but not sold to buyer $i$ with price vector $(p^{*}_{1},\cdots,p^{*}_{n})$. Since $p^{\prime}_{i}\leq p^{*}_{i}$ holds for all $i$, this is the only way to incur a loss when running $(p^{\prime}_{1},\cdots,p^{\prime}_{n})$. Besides, notice that $\sum_{i\in[n]}\mbox{\rm\bf Pr}\left[E_{i}\right]\leq 1$, because at most one event $E_{i}$ can happen. Therefore, it’s sufficient to show that whenever $E_{i}$ happens, the expected loss is bounded by $\epsilon$.
Let $\mathsf{Opt}(S_{i})$ be the expected revenue from the buyers $i,i+1,\cdots,n$ when playing prices $(p^{*}_{1},\cdots,p^{*}_{n})$. When $E_{i}$ happens, we get expected revenue $\mathsf{Opt}(S_{i+1})$ when playing $(p^{*}_{1},\cdots,p^{*}_{n})$ and get $p^{\prime}_{i}$ when playing $(p^{\prime}_{1},\cdots,p^{\prime}_{n})$. Notice that we have
$\mathsf{Opt}(S_{i+1})-p^{\prime}_{i}~{}\leq~{}p^{*}_{i}-p^{\prime}_{i}~{}\leq~{}\epsilon,$
where the first inequality is from the fact that $p^{*}_{i}\geq\mathsf{Opt}(S_{i+1})$, because
$R(1,\cdots,1,p,p^{*}_{i+1},\cdots,p^{*}_{n})~{}<~{}\mathsf{Opt}(S_{i+1})~{}=~{}R(1,\cdots,1,\mathsf{Opt}(S_{i+1}),p^{*}_{i+1},\cdots,p^{*}_{n})$
holds for any $p<\mathsf{Opt}(S_{i+1})$, i.e., $p$ can’t be the maximizer of $R(1,\cdots,1,p,p^{*}_{i+1},\cdots,p^{*}_{n})$ when $p<\mathsf{Opt}(S_{i+1})$. Therefore, when $E_{i}$ happens, the expected loss is bounded by $\epsilon$. ∎
4.2 Algorithm for Discrete Buyers
Now we give our algorithm for buyers with discrete support. Our main goal is to prove the following 4.3:
Theorem 4.3.
For BSPP problem with $n$ buyers having valuations in $V=\{v_{1},v_{2},\cdots,v_{k}\}\subseteq[0,1]$, there exists an $O(n^{2.5}\sqrt{kT}\log T+n^{5}k\log^{2}T)$ regret algorithm.
With the help of 4.3, combining 4.2 with 4.3 proves the main4.1. Therefore, proving 4.3 is sufficient to give the desired $T^{2/3}$ regret bound for general buyers.
Proof of 4.1.
We take $k=n^{-5/3}T^{1/3}$ and uniformly discretize $[0,1]$ into $k$ prices. 4.2 suggests that this leads to an extra $\frac{1}{k}=n^{5/3}T^{-1/3}$ error, so the total regret from the discretization is $T\cdot\frac{1}{k}=O(n^{5/3}T^{2/3})$. On the other hand, the total regret from the algorithm given by 4.3 is $O(n^{2.5}\sqrt{kT}\log T+n^{5}k\log^{2}T)=O(n^{5/3}T^{2/3}\log T+n^{10/3}T^{1/3}\log^{2}T)$. In this paper, we further assume that $n^{5/3}\log T=o(T^{1/3})$, i.e., $n^{5}\log^{3}T=o(T)$, otherwise an $\widetilde{O}(\mathsf{poly}(n))$ algorithm is trivial. Then, the total regret is $O(n^{5/3}T^{2/3}\log T)$. ∎
4.2.1 Notation
Our proof of 4.3 starts from introducing the notations. We use the same set of the notations as the regular distributions algorithm, which are restated for convenience:
  • $\delta$: We define $\delta=\frac{\epsilon}{100n^{2}}$ to be the scaled error parameter. $\delta$ also represents the unit length of those confidence intervals given by concentration inequalities.
  • $R(p_{1},\cdots,p_{n})$ and $p^{*}_{1},\cdots,p^{*}_{n}$: $R(p_{1},\cdots,p_{n})$ represents the expected revenue with prices $(p_{1},\cdots,p_{n})$, and $(p^{*}_{1},\cdots,p^{*}_{n})$ represents the optimal prices that maximizes $R(p_{1},\cdots,p_{n})$.
  • $(\hat{p}_{1},\cdots,\hat{p}_{n})$: The approximately optimal price vector we find in first-step algorithm.
  • $S_{i}$: The suffix buyers $i,i+1,\cdots,n$. When finding $\hat{p}_{i-1}$, the performance of buyers $i,i+1,\cdots,n$ can be seen as a whole. $S_{i}$ represents the whole of this suffix starting from buyer $i$.
  • $P_{i}$: We define $P_{i}~{}:=~{}\prod_{j<i}F_{j}(r_{j}),$ to be the maximum probability of testing buyer $i$.
  • $\mathsf{Pre}(S_{i})$: We define $\mathsf{Pre}(S_{i}):=R(r_{1},\cdots,r_{i-1},1,\cdots,1)$ to be the expected revenue from the buyers before $S_{i}$ with prices $r_{1},\cdots,r_{i-1}$.
  • $\mathsf{Opt}(S_{i})$: We define $\mathsf{Opt}(S_{i}):=R(1,\cdots,1,p^{*}_{i},\cdots,p^{*}_{n})$ to be the optimal revenue from $S_{i}$.
  • $\mathsf{Rev}(S_{i})$: We define $\mathsf{Rev}(S_{i}):=R(1,\cdots,1,\hat{p}_{i},\cdots,\hat{p}_{n})$ to be the near-optimal revenue from $S_{i}$ with near-optimal prices $\hat{p}_{i},\cdots,\hat{p}_{n}$.
  • $\mathsf{Loss}(S_{i})$: We define $\mathsf{Loss}(S_{i}):=\mathsf{Opt}(S_{i})-\mathsf{Rev}(S_{i})$, i.e., $\mathsf{Loss}(S_{i})$ represents the loss from $S_{i}$ when playing prices $\hat{p}_{i},\cdots,\hat{p}_{n}$.
  • $R_{i}(p)$: We define $R_{i}(p)~{}:=~{}R\big{(}r_{1},\cdot,r_{i-1},p_{i}=p,\hat{p}_{i+1},\cdots,\hat{p}_{n}\big{)}.$ It should be noted that $R_{i}(p)$ can be also written as $R_{i}(p)~{}=~{}\mathsf{Pre}(S_{i})+P_{i}\big{(}p\cdot(1-F_{i}(p))+\mathsf{Rev}(S_{i+1})\cdot F_{i}(p)\big{)}.$
  • $R^{*}_{i}(p)$: We define $R^{*}_{i}(p)~{}:=~{}R(r^{s}_{1},\cdots,r^{s}_{i-1},p_{i}=p,p^{*}_{i+1},\cdots,p^{*}_{n})$. It should be noted that $R^{*}_{i}(p)$ can be also written as
    $R^{*}_{i}(p)~{}=~{}\mathsf{Pre}(S_{i})+P_{i}(p\cdot(1-F_{i}(p))+\mathsf{Opt}(S_{i+1})\cdot F_{i}(p))~{}=~{}R_{i}(p)+P_{i}\cdot F_{i}(p)\cdot\mathsf{Loss}(S_{i+1}).$
We also introduce the following notations for the discretized setting:
  • $V=\{v_{1},\cdots,v_{k}\}$: The discrete values of buyers. Every buyer’s value must fall in $V$.
  • $\mathsf{CP}_{i}$: The candidate prices set of buyer $i$. Initially $\mathsf{CP}_{i}$ is set to be $V$.
  • $r_{i}$: we define $r_{i}=\max\mathsf{CP}_{i}$ to be the maximum feasible price for buyer $i$.
4.2.2 High-Level Structure: Shrinking Candidate Prices Sets
The high-level structure of our algorithm is to gradually shrink the sets of candidate prices. Specifically, we first give the following main sub-routine:
Lemma 4.4.
Given a BSPP instance with $R(p_{1},\cdots,p_{n})$ being the expected revenue function and $(p^{*}_{1},\cdots,p^{*}_{n})$ being the optimal prices. Assume the valuation of each single buyer falls in the set $V=\{v_{1},\cdots,v_{k}\}$. Given error parameter $\epsilon\geq 1/\sqrt{T}$ and candidate sets $\mathsf{CP}_{1}\subseteq V,\cdots,\mathsf{CP}_{n}\subseteq V$ satisfying $p^{*}_{i}\in\mathsf{CP}_{i}$ for all $i\in[n]$, there exists an algorithm that tests $O(\frac{n^{5}k\log T}{\epsilon^{2}})$ rounds, where each tested price vector $(p_{1},\cdots,p_{n})$ satisfies $p_{i}\in\mathsf{CP}_{i}$ for all $i\in[n]$, and outputs with probability $1-T^{-3}$ new candidate sets $\mathsf{CP}^{\prime}_{1}\subseteq\mathsf{CP}_{1},\cdots,\mathsf{CP}^{\prime}_{n}\subseteq\mathsf{CP}_{n}$ satisfying
  • For $i\in[n]$, $p^{*}_{i}\in\mathsf{CP}^{\prime}_{i}$.
  • For every price vector $(p_{1},\cdots,p_{n})$ that satisfies $p_{i}\in\mathsf{CP}^{\prime}_{i}$ for all $i\in[n]$, we have
    $R(p^{*}_{1},\cdots,p^{*}_{n})-R(p_{1},\cdots,p_{n})\leq\epsilon.$
We defer the detailed algorithm and proofs for 4.4 to Section 4.2.3 and first give the proof for 4.3. The main idea is to use 4.4 for $O(\log T)$ times and gradually halve the one-round error when playing with prices inside the candidate sets.
Input: Hidden Revenue Function $R(p_{1},\cdots,p_{n})$, time horizon $T$
1 Let $\epsilon_{1}\leftarrow 1,\mathsf{CP}^{(1)}_{j}\leftarrow V=\{v_{1},\cdots,v_{k}\}$ for all $j\in[n]$, and $i\leftarrow 1$.
2 while $\epsilon_{i}>\frac{n^{2.5}\sqrt{k}\log T}{\sqrt{T}}$ do
3       Run the algorithm described in 4.4 with $\epsilon_{i},\mathsf{CP}^{(i)}_{1},\cdots,\mathsf{CP}^{(i)}_{n}$ as input, and get $\mathsf{CP}^{\prime}_{1},\cdots,\mathsf{CP}^{\prime}_{n}$. Let $\mathsf{CP}^{(i+1)}_{j}\leftarrow\mathsf{CP}^{\prime}_{j}$ for all $j\in[n]$.
4       Let $\epsilon_{i+1}\leftarrow\frac{1}{2}\epsilon_{i}$ and $i\leftarrow i+1$.
5
Finish remaining rounds with any $(p_{1},\cdots,p_{n})$ satisfying $p_{j}\in\mathsf{CP}_{j}$ for all $j\in[n]$.
ALGORITHM 9 $O(n^{2.5}\sqrt{kT}\log T)$ Regret Algorithm
Proof of 4.3.
We claim that the regret of Algorithm 9 is $O(n^{2.5}\sqrt{kT}\log T)$ with probability $1-T^{-2}$. Algorithm 9 uses 4.4 for multiple times with a halving error parameter. Assume Algorithm 9 ends with $i=q+1$, i.e., the while loop runs $q$ times. Since we obtain $\epsilon_{i+1}=\epsilon_{i}/2$ and $\epsilon_{q}>\frac{n^{2.5}\sqrt{k}\log T}{\sqrt{T}}$, there must be $q=O(\log T)$.
For $i\in[q]$, let $\mathsf{Alg}_{i}$ represent the corresponding algorithm we call when using 4.4 with $\epsilon_{i},\{\mathsf{CP}^{(i)}_{j}\}$ as the input. To use 4.4 we need to verify that $p^{*}_{j}\in\mathsf{CP}^{(i)}_{j}$ for all $i\in[q],j\in[n]$. The condition $p^{*}_{j}\in\mathsf{CP}^{(1)}_{j}=V$ holds because every buyer only has valuation in $V$. For $i=2,3,\cdots,k$, the condition $p^{*}_{j}\in\mathsf{CP}^{(i)}_{j}$ is guaranteed by 4.4 when calling $\mathsf{Alg}_{i-1}$. The failing probability of 4.4 is $T^{-3}$. By the union bound, with probability $1-T^{-2}$, 4.4 always holds.
Now we give the regret of Algorithm 9. 4.4 guarantees that for $i\geq 2$, every price vector we test in $\mathsf{Alg}_{i}$ has regret bounded by $\epsilon_{i-1}$. Therefore, the total regret of Algorithm 9 can be bounded by
$1\cdot O\big{(}\epsilon_{1}^{-2}n^{5}k\log^{2}T\big{)}~{}+~{}\sum_{i=2}^{q}\epsilon_{i-1}\cdot O\big{(}\epsilon^{-2}_{i}n^{5}k\log^{2}T\big{)}~{}+~{}T\cdot\epsilon_{q}~{}=~{}O(n^{5}k\log^{2}T+n^{2.5}\sqrt{kT}\log T).$
Finally, we also need to check that Algorithm 9 uses no more than $T$ rounds. If $n^{5}k\log^{2}T>T$, this implies that $\epsilon_{1}<\frac{n^{2.5}\sqrt{k}\log T}{\sqrt{T}}$, so the while loop will not be processed. Otherwise, 4.4 suggests that $\mathsf{Alg}_{i}$ runs $O(\epsilon^{-2}_{i}n^{5}k\log^{2}T)$ rounds. So the total number of rounds in the while loop is $\sum_{i\in[k]}O(\epsilon^{-2}_{i}n^{5}k\log^{2}T)~{}=~{}O(T).$ Therefore, Algorithm 9 is feasible. ∎
4.2.3 Main Sub-routine: Proof of 4.4
We first present our main sub-routine to prove 4.4. Similar to the regular distributions setting, there are two major steps in the sub-routine: we first find $(\hat{p}_{1},\cdots,\hat{p}_{n})$ that approximates $(p^{*}_{1},\cdots,p^{*}_{n})$, and the second step is to use the near-optimal prices as a benchmark to update the candidate price sets. Specifically, we have the following two separate lemmas for these two steps:
Lemma 4.5.
Given $\epsilon\geq\frac{1}{\sqrt{T}}$, candidate sets $\mathsf{CP}_{1},\cdots,\mathsf{CP}_{n}$ satisfying $p^{*}_{i}\in\mathsf{CP}_{i}$, there exists an algorithm that tests $O(\frac{n^{5}k\log T}{\epsilon^{2}})$ rounds with price vectors $(p_{1},\cdots,p_{n})$ that satisfy $p_{i}\in\mathsf{CP}_{i}$ for all $i\in[n]$ and outputs $(\hat{p}_{1},\cdots,\hat{p}_{n})$. With probability $1-T^{-4}$, we have $R_{i}(p^{*}_{i})-R_{i}(\hat{p}_{i})\leq 2\delta$ for all $i\in[n]$, where $\delta:=\frac{\epsilon}{100n^{2}}$ is the scaled error parameter.
Lemma 4.6.
Given $\epsilon\geq\frac{1}{\sqrt{T}}$, candidate sets $\mathsf{CP}_{1},\cdots,\mathsf{CP}_{n}$ satisfying $p^{*}_{i}\in\mathsf{CP}_{i}$, and $(\hat{p}_{1},\cdots,\hat{p}_{n})$ that satisfies $R_{i}(p^{*}_{i})-R_{i}(\hat{p}_{i})\geq 2\delta$, there exists an algorithm that tests $O(\frac{n^{5}k\log T}{\epsilon^{2}})$ rounds with price vectors $(p_{1},\cdots,p_{n})$ that satisfy $p_{i}\in\mathsf{CP}_{i}$ for all $i\in[n]$ and outputs with probability $1-T^{-4}$ $\mathsf{CP}^{\prime}_{1}\subseteq\mathsf{CP}_{1},\cdots,\mathsf{CP}^{\prime}_{n}\subseteq\mathsf{CP}_{n}$ that satisfies:
  • $p^{*}_{i}\in\mathsf{CP}^{\prime}_{i}$.
  • For $p_{1},\cdots,p_{n}$ that satisfies $p_{i}\in\mathsf{CP}^{\prime}_{i}$ for $i\in[n]$, we have $R(p^{*}_{1},\cdots,p^{*}_{n})-R(p_{1},\cdots,p_{n})\leq\epsilon$.
Combining 4.5 and 4.6 with a union bound directly proves 4.4. It only remains to prove the two lemmas. Now we prove them separately in the following subsections.
4.2.4 Proof of 4.5: Finding Approximately Good Prices
The key idea of 4.5 is simply enumerating all prices in the candidate set and picking the best one during the test. Specifically, Algorithm 10 is the desired algorithm for 4.5.
Input: Hidden function $R(p_{1},\cdots,p_{n})$, candidate sets $\mathsf{CP}_{1},\cdots,\mathsf{CP}_{n}$, scaled error parameter $\delta$
1 For $i\in[n]$, let $r_{i}$ be the largest price in $\mathsf{CP}_{i}$.
2 for $i=n\to 1$ do
3       For $j<i$, fix $p_{j}=r_{j}$; for $j>i$, fix $p_{j}=\hat{p}_{j}$.
4       for $p\in\mathsf{CP}_{i}$ do
5             Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=p$. Let $\hat{R}_{i}(p)$ be the average reward.
6      Let $\hat{p}_{i}=\arg\max_{p\in\mathsf{CP}_{i}}\hat{R}_{i}(p)$.
Output: $\hat{p}_{i}$
ALGORITHM 10 Finding Approximately Good Prices for General Distributions
Proof of 4.5.
For the number of rounds of Algorithm 10, our algorithm enumerates all prices in the candidate sets, so there are $O(nk)$ prices. Each price is tested for $O(\frac{\log T}{\delta^{2}})=O(\frac{n^{4}\log T}{\epsilon^{2}})$ rounds, so Algorithm 10 tests $O(\frac{n^{5}k\log T}{\epsilon^{2}})$ rounds. It remains to show that $R_{i}(\hat{p}_{i})\geq R_{i}(p^{*}_{i})-2\delta$. We start from following claim:
Claim 4.7.
In Algorithm 10, $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ simultaneously holds for every $i\in[n]$ and $p\in\mathsf{CP}_{i}$ with probability at least $1-T^{-4}$.
Proof.
In Algorithm 10, $\hat{R}_{i}(p)$ is estimated via $N=C\cdot\frac{\log T}{\delta^{2}}$ samples. By Hoeffding’s Inequality,
$\mbox{\rm\bf Pr}\left[|\hat{R}_{i}(p)-R_{i}(p)|>\delta\right]~{}\leq~{}2\exp(-2N\delta^{2})~{}=~{}2T^{-2C}~{}<~{}T^{-5}.$
The last inequality holds when $C$ is a constant greater than $3$. Then, we have $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ holds with probability $1-T^{-5}$ for a single tested price $p$.
Algorithm 10 tests $\sum_{i}|\mathsf{CP}_{i}|\leq nk$ prices. In this paper, we assume that $n\cdot k\leq T$, otherwise it’s not reasonable for not being able to test every price in $T$ rounds. Therefore, by the union bound, we have $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ holds for all tested price $p$ with probability $1-T^{-4}$. ∎
With 4.7, it is straight forward to show $\hat{p}_{i}$ is approximately optimal: we have
$R_{i}(\hat{p}_{i})~{}\geq~{}\hat{R}_{i}(\hat{p}_{i})-\delta~{}\geq~{}\hat{R}_{i}(p^{*}_{i})-\delta~{}\geq~{}R_{i}(p^{*}_{i})-2\delta.$
Here we use 4.7 in the first and the third inequality, and the second inequality is from the fact that $\hat{p}_{i}$ maximizes $\hat{R}_{i}(p)$. ∎
4.2.5 Proof of 4.6: Shrinking Candidate Prices Sets
Before proving 4.6, we first introduce the following claim, which shows the loss of playing prices $\hat{p}_{i},\cdots,\hat{p}_{n}$ for the suffix $S_{i}$ is small:
Claim 4.8.
Given $(\hat{p}_{1},\cdots,\hat{p}_{n})$ that satisfies $R_{i}(p^{*}_{i})-R_{i}(\hat{p}_{i})\leq 2\delta$ for $i\in[n]$. Recall that $\mathsf{Loss}_{i}:=R(1,\cdots,1,p^{*}_{i},\cdots,p^{*}_{n})-R(1,\cdots,1,\hat{p}_{i},\cdots,\hat{p}_{n})$. We have $P_{i}\cdot\mathsf{Loss}_{i}\leq 2(n-i+1)\delta$.
Proof.
We prove 4.8 via induction. The base case $i=n$ is true because of the assumption $P_{n}\cdot\mathsf{Loss}_{n}=R_{n}(p^{*}_{n})-R_{n}(\hat{p}_{n})\leq 2\delta$.
For the induction step, we will show $P_{i}\cdot\mathsf{Loss}_{i}\leq 2(n-i+1)\delta$ under the induction hypothesis $P_{i+1}\mathsf{Loss}_{i+1}\leq 2(n-i)\delta$. We have
$P_{i}\cdot\mathsf{Loss}(S_{i})~{}=~{}$ $R(r_{1},\cdots,r_{i-1},p^{*}_{i},p^{*}_{i+1}\cdots,p^{*}_{n})-R(r_{1},\cdots,r_{i-1},\hat{p}_{i},\hat{p}_{i+1}\cdots,\hat{p}_{n})$
[…]
$+R(r_{1},\cdots,r_{i-1},p^{*}_{i},\hat{p}_{i+1}\cdots,\hat{p}_{n})-R(r_{1},\cdots,r_{i-1},\hat{p}_{i},\hat{p}_{i+1}\cdots,\hat{p}_{n})$
[…]
Here, the first inequality uses the fact that $F_{i}(p^{*}_{i})\cdot P_{i}\leq F_{i}(r_{i})\cdot P_{i}=P_{i+1}$, and the second inequality uses the assumption that $P_{i+1}\mathsf{Loss}_{i+1}\leq 2(n-i)\delta$, and $R_{i}(p^{*}_{i})-R_{i}(\hat{p}_{i})\leq 2\delta$. ∎
Next, similar to the regular distributions setting, we use $R^{*}_{i}(p):=R(r_{1},\cdots,r_{i-1},p,p^{*}_{i+1},\cdots,p^{*}_{n})$ as a bridge. With the help of 4.8, we have the following claim:
Claim 4.9.
Given $(\hat{p}_{1},\cdots,\hat{p}_{n})$ that satisfies $R_{i}(p^{*}_{i})-R_{i}(\hat{p}_{i})\leq 2\delta$ for $i\in[n]$. We have $R_{i}(p)\leq R^{*}_{i}(p)\leq R_{i}(p)+2(n-i)\delta$ for all $i\in[n]$.
Proof.
The first inequality $R_{i}(p)\leq R^{*}_{i}(p)$ holds because of the optimality of prices $p^{*}_{i+1},\cdots,p^{*}_{n}$. The second inequality $R^{*}_{i}(p)\leq R_{i}(p)+2(n-i+1)\delta$ holds from the fact that $R^{*}_{i}(p)-R_{i}(p)=F_{i}(p)\cdot P_{i}\cdot\mathsf{Loss}_{i+1}\leq P_{i+1}\cdot\mathsf{Loss}_{i+1}\leq 2(n-i)\delta$, where the last inequality is from 4.8. ∎
The key idea of 4.6 is to keep every price which is possible to become the maximzer of $R^{*}_{i}(p)$. However, directly estimating $R^{*}_{i}(p)$ is not possible. Therefore, we first estimate $R_{i}(p)$, and use 4.9 to bound the value of $R^{*}_{i}(p)$. Specifically, Algorithm 11 is the desired algorithm for 4.6.
Input: Hidden function $R(p_{1},\cdots,p_{n})$, candidate sets $\mathsf{CP}_{1},\cdots,\mathsf{CP}_{n}$, scaled error parameter $\delta$, and approximately good prices $(\hat{p}_{1},\cdots,\hat{p}_{n})$ from Algorithm 10
1 For $i\in[n]$, let $r_{i}$ be the largest price in $\mathsf{CP}_{i}$.
2 For $i\in[n]$, initiate $\mathsf{CP}^{\prime}_{i}=\mathsf{CP}_{i}$. for $i=n\to 1$ do
3       For $j<i$, fix $p_{j}=r_{j}$; for $j>i$, fix $p_{j}=\hat{p}_{j}$.
4       Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=\hat{p}_{i}$. Let $\hat{R}_{i}(\hat{p}_{i})$ be the average reward. for $p\in\mathsf{CP}^{\prime}_{i}\setminus\{\hat{p}_{i}\}$ do
5             Test $C\cdot\frac{\log T}{\delta^{2}}$ rounds with $p_{i}=p$. Let $\hat{R}_{i}(p)$ be the average reward.
6             Remove $p$ from $\mathsf{CP}^{\prime}_{i}$ if $\hat{R}_{i}(p)<\hat{R}_{i}(\hat{p}_{i})-2(n-i+1)\delta-2\delta$.
7
Output: Updated sets $\mathsf{CP}^{\prime}_{1},\cdots,\mathsf{CP}^{\prime}_{n}$
ALGORITHM 11 Shrinking Candidate Sets via Approximately Good Prices
Proof of 4.6.
For the number of rounds of Algorithm 10, our algorithm enumerates all prices in the candidate sets, so there are $O(nk)$ prices. Each price is tested for $O(\frac{\log T}{\delta^{2}})=O(\frac{n^{4}\log T}{\epsilon^{2}})$ rounds, so Algorithm 10 tests $O(\frac{n^{5}k\log T}{\epsilon^{2}})$ rounds. It remains to show $p^{*}_{i}\in\mathsf{CP}^{\prime}_{i}$ and give the loss bound for playing prices in $\mathsf{CP}^{\prime}_{1}\cdots,\mathsf{CP}^{\prime}_{n}$. We start from the following claim:
Claim 4.10.
In Algorithm 11, $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ simultaneously holds for every $i$ and $p$ with probability at least $1-T^{-4}$.
Proof.
In Algorithm 11, $\hat{R}_{i}(p)$ is estimated via $N=C\cdot\frac{\log T}{\delta^{2}}$ samples. By Hoeffding’s Inequality,
$\mbox{\rm\bf Pr}\left[|\hat{R}_{i}(p)-R_{i}(p)|>\delta\right]~{}\leq~{}2\exp(-2N\delta^{2})~{}=~{}2T^{-2C}~{}<~{}T^{-5}.$
The last inequality holds when $C$ is a constant greater than $3$. Then, we have $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ holds with probability $1-T^{-5}$ for a single tested price $p$.
Algorithm 11 tests $\sum_{i}|\mathsf{CP}_{i}|\leq nk$ prices. In this paper, we assume that $n\cdot k\leq T$, otherwise it is not reasonable for not being able to test every price in $T$ rounds. Therefore, by the union bound, we have $|\hat{R}_{i}(p)-R_{i}(p)|\leq\delta$ holds for all tested price $p$ with probability $1-T^{-4}$. ∎
With 4.10, it is sufficient to show that a price $p$ is not possible to become $p^{*}_{i}$ if $\hat{R}_{i}(p)<\hat{R}_{i}(\hat{p}_{i})-2(n-i+1)\delta-2\delta$ holds: Notice that
$\hat{R}_{i}(p^{*}_{i})~{}$ $\geq~{}R_{i}(p^{*}_{i})-\delta$
$\geq~{}R^{*}_{i}(p^{*}_{i})-2(n-i+1)\delta-\delta$
$\geq~{}R^{*}_{i}(\hat{p}_{i})-2(n-i+1)\delta-\delta$
$\geq~{}R_{i}(\hat{p}_{i})-2(n-i+1)\delta-\delta$
$\geq~{}\hat{R}_{i}(\hat{p}_{i})-2(n-i+1)\delta-2\delta.$
Here, the first inequality and the last inequality is from 4.10. The second and the fourth inequality is from 4.9. The third inequality is from the fact that $p^{*}_{i}$ maximizes $R^{*}_{i}(p)$. Therefore, if a price satisfies $\hat{R}_{i}(p)<\hat{R}_{i}(\hat{p}_{i})-2(n-i+1)\delta-2\delta$, this price $p$ can’t be a candidate of $p^{*}_{i}$.
After Algorithm 11, every price $p$ in $\mathsf{CP}^{\prime}_{i}$ must satisfy $\hat{R}_{i}(p)\geq\hat{R}_{i}(\hat{p}_{i})-2(n-i)\delta-4\delta$. Then, 4.10 and the assumption $R_{i}(\hat{p}_{i})\geq R_{i}(p^{*}_{i})-2\delta$ gives
$R_{i}(p)~{}\geq~{}R_{i}(p)-\delta~{}\geq~{}R_{i}(\hat{p}_{i})-2(n-i)\delta-6\delta~{}\geq~{}R_{i}(p^{*}_{i})-2(n-i)\delta-8\delta.$
Combining the above inequality with 4.9, we have
$R^{*}_{i}(p)~{}\geq~{}R_{i}(p)~{}\geq~{}R_{i}(p^{*}_{i})-2(n-i)\delta-8\delta~{}\geq~{}R^{*}_{i}(p^{*}_{i})-4(n-i)\delta-10\delta.$
Now we show that this is sufficient to bound the loss of playing prices in sets $\mathsf{CP}^{\prime}_{1},\cdots,\mathsf{CP}^{\prime}_{n}$. Specifically, we will prove the following statement via induction: Given $p_{1},p_{2},\cdots,p_{i}$, such that $p_{j}\in\mathsf{CP}^{\prime}_{j}$ for $j\in[i]$, we have
$R(p^{*}_{1},\cdots,p^{*}_{i},p^{*}_{i+1},\cdots,p^{*}_{n})-R(p_{1},\cdots,p_{i},p^{*}_{i+1},\cdots,p^{*}_{n})~{}\leq~{}20in\cdot\delta.$
The base case is $i=1$. The statement holds because we have
$R(p^{*}_{1},p^{*}_{2},\cdots,p^{*}_{n})-R(p_{1},p^{*}_{2},\cdots,p^{*}_{n})~{}=~{}R_{1}(p^{*}_{1})-R_{1}(p_{1})~{}\leq~{}(4n+6)\delta\leq 20n\delta.$
For the induction step, we have
$R(p^{*}_{1},\cdots,p^{*}_{i-1},p^{*}_{i},p^{*}_{i+1},\cdots,p^{*}_{n})-R(p_{1},\cdots,p_{i-1},p_{i},p^{*}_{i+1},\cdots,p^{*}_{n})$
[…]
$+R(p_{1},\cdots,p_{i-1},p^{*}_{i},p^{*}_{i+1},\cdots,p^{*}_{n})-R(p_{1},\cdots,p_{i-1},p_{i},p^{*}_{i+1},\cdots,p^{*}_{n}).$
[…]
Here, the first inequality is from the induction hypothesis. The second inequality is from the fact that $p_{j}\leq r_{j}$, and therefore $\prod_{j<i}F_{j}(p_{j})\leq\prod_{j<i}F_{j}(r_{j})=P_{i}$. It should also be noted that the inequality can be applied only when we have $R^{*}_{i}(p^{*}_{i})\geq R^{*}_{i}(p_{i})$, which is true because $p^{*}_{i}$ maximizes $R^{*}_{i}(p)$. The last inequality is from the fact that $R^{*}_{i}(p)\geq R^{*}_{i}(p^{*}_{i})-4(n-i)\delta-10\delta\geq R^{*}_{i}(p^{*}_{i})-20n\delta$.
Finally, taking $i=n$ in the above statement together with the fact that $20n^{2}\delta\leq\epsilon$ finishes the proof of 4.6. ∎
5 Linear Regret Lower Bound for Adversarial Valuations
We show an $\Omega(T)$ regret lower bound for learning sequential posted pricing with adversarial buyer values. [KL03] gave an $\widetilde{O}(T^{{2}/{3}})$ regret upper bound for bandit sequential posted pricing with adversarial values when there is only a single buyer. The following result shows that it is not possible to generalize their result to multiple buyers.
Theorem 5.1.
For Online Sequential Posted Pricing problem with oblivious adversarial inputs, there exists an instance with $n=2$ buyers such that the optimal fixed-threshold strategy has total revenue $\frac{3}{4}T$ but no online algorithm can obtain total value more than $\frac{1}{2}T$.
Proof.
The proof follows the hardness example for Online Prophet Inequality problem in [GKSW24]. Here we restate the example for completeness.
Let $s$ be a binary string in $\{0,1\}^{T}$. Define $Bin(s)$ to be the binary decimal corresponding to $s$. For example, $Bin(1000)=(0.1000)_{2}=\frac{1}{2}$ and $Bin(0101)=(0.0101)_{2}=\frac{5}{16}$.
At the beginning, the adversary chooses a $T$-bits string $s=s_{1}s_{2},\cdots,s_{T}$ uniformly at random, i.e., $s_{i}$ is set to be 00 or $1$ independently with probability $\frac{1}{2}$. In the $i$-th round, the value of the first buyer will be $v_{1}=\frac{1}{2}+\varepsilon\cdot\alpha_{i}$, where $\alpha_{i}$ is set to be $Bin(s_{1}s_{2}...s_{i-1}+0+1^{T-i+1})$, and $\varepsilon$ is an arbitrarily small constant that doesn’t effect the value. The value of the second buyer only depends on $s_{i}$: $v_{2}$ is set to be 00 when $s_{i}=1$, while $v_{i}=1$ when $s_{i}=0$.
The key idea of this example is that we have $Bin(s)>\alpha_{i}$ when $s_{i}$ is 00, and $Bin(s)<\alpha_{i}$ when $s_{i}=1$. Therefore, if we set $p_{1}=Bin(s)$ and $p_{2}=1$, we can receive $v_{2}$ when $v_{2}$ is $1$, and otherwise $v_{1}$. Since $s$ is generated uniformly at random, the expected revenue is $\frac{3}{4}T$. However, for any online algorithm, it only knows that the value of $v_{2}$ is 00 or $1$ with probability $\frac{1}{2}$. Therefore, it can only get revenue $\frac{1}{2}$ in expectation and the maximum total revenue is $\frac{1}{2}T$. ∎
Acknowledgements
We are thankful to Thomas Kesselheim as the project was initiated in discussions with him. We also thank the anonymous reviewers who helped greatly improve the presentation of the paper.